### v2

# Ikeda type lift on ${\rm SO}(3,n+1)$

Abstract.
By using Ikeda’s theory for a compatible family of Eisenstein series, we explicitly construct Ikeda type lifts on the special orthogonal group $G={\rm SO}(3,n+1)$ over $\mathbb{Q}$ with $n\geq 3$ which splits everywhere at finite places. Our lifts are Hecke eigen cusp forms of weight $l$ ($l\geq n+2$, even) and come from elliptic newforms with respect to ${\rm SL}_{2}(\mathbb{Z})$ which are of weight $l-\frac{n-2}{2}$ when $n$ is even and 2ln+12l-n+1 when $n$ is odd.
Key words and phrases:
Ikeda type lifts, Eisenstein series, Siegel series, Langlands functoriality
2010 Mathematics Subject Classification:
Primary 11F55, Secondary 11F70, 22E55
The first author is partially supported by NSERC grant #482564.
1. Introduction
Constructing cusp forms on reductive groups is an important subject in both the theory of automorphic representations and number theory. There are several important works involving trace formulas [1], theta lifts [25] (see [20] as a related work), the descend method [6] or converse theorems [7], but in this paper, we focus on Ikeda type construction (see [10], [11], [14], [15],[16]).
Let $\mathbb{A}=\mathbb{A}_{\mathbb{Q}}$ be the ring of adeles of $\mathbb{Q}$. For an integer $n\geq 3$, let $A$ be a positive definite integral matrix of size n2n-2. Assume $A$ is even (namely, any diagonal entries are even integers) if $n$ is even while $A={\rm diag}(1,A^{\prime})$ if $n$ is odd where $A^{\prime}$ is a positive definite integral matrix of size n3n-3. Let $J_{2,n}={\rm antidiag}(1,1,-A,1,1)$ and $J_{3,n+1}={\rm antidiag}(1,1,1,-A,1,1,1)$ which are of size $n+3$ and $n+4$ respectively (see Section 2). Let $G={\rm SO}(3,n+1)$ be the special orthogonal group over $\mathbb{Q}$ associated to the symmetric pairing defined by $J_{3,n+1}$. Note that if $n$ is even, $G$ does not have discrete series. Assume $G$ splits everywhere at finite places. By [26, Section 2.1], it is equivalent to $\det(A)=1$ and $n\equiv 2$ mod 8 when $n$ is even, $\det(A^{\prime})=1$ and $n\equiv 3$ mod 8 when $n$ is odd. For instance, this condition is satisfied when $n=8a+2$, $A={\rm diag}(\overbrace{A_{8},\ldots,A_{8}}^{a})$ or $n=8a+3$, $A={\rm diag}(1,\overbrace{A_{8},\ldots,A_{8}}^{a})$ where $A_{8}$ is the $E_{8}$ Cartan matrix given as an element of even integral matrix in $M_{8}(\mathbb{Z})$. This agrees with the result of [8] since $\frac{n+4}{2}\equiv 3$ (mod 4) if $n$ is even, and $\frac{n+3}{2}\equiv 3$ (mod 4).
Let $P$ be the Siegel parabolic subgroup of $G$ with the Levi decomposition $P=MN$ where $M=\{{\rm diag}(t,m,t^{-1})\ |\ t\in{\rm GL}_{1},\ m\in{\rm SO}(2,n)\}\simeq{\rm GL}_{1}\times{\rm SO}(2,n)$ and $N\simeq\mathbb{G}^{n+2}_{a}$ where ${\rm SO}(2,n)$ is defined by $J_{2,n}$. Let $V^{\prime}$ be the quadratic space associated to $J_{2,n}$ with the quadratic map $q:V^{\prime}\longrightarrow\mathbb{G}_{a}$ and we identify $V^{\prime}$ with $N$. We also denote by $(\ast,\ast)_{V^{\prime}}$ the corresponding symmetric bilinear pairing.
In [23, Section 4], Pollack defined the Eisenstein series $E_{l}$ on $G(\mathbb{A})$ of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$), and computed unramified and archimedean part of Fourier coefficients. He can show the algebraicity of the Fourier coefficients. However, he did not compute the Siegel series explicitly which come from the bad finite places. The computation of the Siegel series is complicated and it is contained in Section 11. One corollary of our explicit computation is that we can show that Fourier coefficients of Eisenstein series have bounded denominator. (See Theorem 4.8 and 4.9.) After a long computation, we can write down the Fourier expansion of the Eisenstein series as follows:
$E_{l}(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)\geq 0}a_{E_{l}}(\eta)\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=\gamma k_{f}g_{\infty}\in G(\mathbb{A})=G(\mathbb{Q})(G(\widehat{\mathbb{Z}})\times G(\mathbb{R}))$
where $\mathcal{W}_{2\pi\eta}$ is Pollack’s spherical function associated to the additive character $e^{2\pi\sqrt{-1}(\eta,\ast)_{V^{\prime}}}$ on $V^{\prime}(\mathbb{R})$ (see [23, Section 3] but we use a slightly different normalization as explained later). We remark that $a_{E_{l}}(\eta)=0$ unless $\eta\in V^{\prime}(\mathbb{Z})$.
When $n$ is even ($n>2$), for each $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$ and each rational prime $p$, there exists a Laurent polynomial $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p})=\widetilde{Q}_{\eta,p}(X_{p})\in\mathbb{Z}[X_{p},X^{-1}_{p}]$ depending on $\eta$ and the unramified Schwartz function $\Phi_{p}={\rm char}_{V^{\prime}(\mathbb{Z}_{p})}\in\mathcal{S}(V^{\prime}(\mathbb{Q}_{p}))$ satisfying $\widetilde{Q}_{\eta,p}(X_{p})=\widetilde{Q}_{\eta,p}(X^{-1}_{p})$ such that
$a_{E_{l}}(\eta)=C_{l,n}q(\eta)^{\frac{l-\frac{n}{2}}{2}}\displaystyle\prod_{p}\widetilde{Q}_{\eta,p}(p^{\frac{l-\frac{n}{2}}{2}})$
where the constant $C_{l,n}$ is given explicitly in Theorem 4.8. We remark that $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p})$ is also defined for any $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$. For any $g_{p}={\rm diag}(t_{p},m_{p},t^{-1}_{p})n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$, we will see later that $\widetilde{Q}_{\eta,g_{p}\Phi_{p}}(X_{p})=|t_{p}|^{\frac{n+2}{2}}_{p}\widetilde{Q}_{g^{-1}_{p}\eta,\Phi_{p}}(X_{p})$. Thus, $\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})$ is also a Laurent polynomial (over $\mathbb{Q}$) satisfying $\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})=\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X^{-1}_{p})$.
Now we are ready to explain our main results.
Theorem 1.1 ($n$ even ($n>2$)).
Keep the notation being as above. Let $l>n+1$ be an even integer. Let $f$ be a normalized Hecke eigen cusp form of weight $k:=l-\frac{n-2}{2}$ with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $f(z)=\displaystyle\sum_{m=1}^{\infty}a_{f}(m)m^{\frac{k-1}{2}}q^{m}$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $A_{f}(\eta)(g_{\mathbb{f}})=q(\eta)^{\frac{k-1}{2}}\displaystyle\prod_{p}\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(\alpha_{p})$ for $g_{\mathbb{f}}=(g_{p})_{p}\in G(\mathbb{A}_{\mathbb{f}})$. Then
$F_{f}(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A_{f}(\eta)(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})$
is a non-zero Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F_{f}}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Further, the standard $L$-function of $\Pi_{F_{f}}$ is given by
$L(s,\Pi_{F_{f}})=L(s,{\rm Sym}^{2}\pi_{f})\prod_{i=-\frac{n}{2}}^{\frac{n}{2}}\zeta(s+i).$
When $n$ is odd, we need a slight modification. Let $l>n+1$ be an even integer. Let $f$ be a normalized Hecke eigen cusp form of weight k:=2ln+1k:=2l-n+1 with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $f(z)=\displaystyle\sum_{m=1}^{\infty}a_{f}(m)m^{\frac{k-1}{2}}q^{m}$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $h(z)=\displaystyle\sum_{m>0\atop\text{$(-1)^{\frac{k}{2}}m\equiv 0,1$ (mod 4)}}c(m)q^{m}\in S_{l-\frac{n}{2}+1}(\Gamma_{0}(4))$ be the one corresponding to $f$ by the Shimura correspondence. Write $q(\eta)=\mathfrak{d}_{\eta}\mathfrak{f}_{\eta}^{2}$ such that $\mathfrak{d}_{\eta}$ is the absolute discriminant of $\mathbb{Q}\left(\sqrt{\epsilon q(\eta)}\right)/\mathbb{Q}$, where $\epsilon=\begin{cases}1,&\text{if $q(\eta)_{1}\equiv 1$ (mod 4)}\\ -1,&\text{if $q(\eta)_{1}\equiv 3$ (mod 4)}\end{cases}$, and $q(\eta)=2^{v_{2}(\eta)}q(\eta)_{1}$, $q(\eta)_{1}$ odd. Let $\chi_{\eta}$ be the primitive Dirichlet character corresponding to $\mathbb{Q}\left(\sqrt{\epsilon q(\eta)}\right)/\mathbb{Q}$. Then for each rational prime $p$, there exists a Laurent polynomial $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p})=\widetilde{Q}_{\eta,p}(X_{p})\in\mathbb{Z}[X_{p},X^{-1}_{p}]$ depending on $\eta$ and $\Phi_{p}={\rm char}_{V^{\prime}(\mathbb{Z}_{p})}$ satisfying $\widetilde{Q}_{\eta,p}(X_{p})=\widetilde{Q}_{\eta,p}(X^{-1}_{p})$ such that
$a_{E}(\eta)=C_{l,n}^{\prime}L(\tfrac{n+1}{2}-l,\chi_{\eta})\mathfrak{f}_{\eta}^{l-\frac{n}{2}}\prod_{p}\widetilde{Q}_{\eta,p}(p^{l-\frac{n}{2}}),$
where $C_{l,n}^{\prime}$ is given in Theorem 4.9.
Theorem 1.2 ($n$ odd ($n\geq 3$)).
Let $A_{f}(\eta)(g_{\mathbb{f}})=c(\mathfrak{d}_{\eta})\mathfrak{f}_{\eta}^{\frac{k-1}{2}}\displaystyle\prod_{p|q(\eta)}\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(\alpha_{p})$ for $g_{\mathbb{f}}=(g_{p})_{p}\in G(\mathbb{A}_{\mathbb{f}})$. Then
$F_{f}(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A_{f}(\eta)(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})$
is a non-zero Hecke eigen cusp form of weight $l$ with respect to G(^)G($\widehat{\mathbb{Z}}$). Let $\Pi_{F}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F_{f}$. Further, the standard $L$-function of $\Pi_{F_{f}}$ is given by
$L(s,\Pi_{F_{f}})=L(s,\pi_{f})\prod_{i=0}^{n}\zeta(s+\frac{n}{2}-i).$
We organize this paper as follows. In Section 2, we give explicit forms of $P$ and another maximal parabolic $Q$ which gives rise to the Jacobi group $J$ in our setting. In Section 3, we recall Pollack’s definition of certain types of automorphic forms on $G(\mathbb{A})$ [23]. In Section 4, we recall the definition of Eisenstein series, and by assuming the explicit calculation of the Siegel series in Section 11, we write down the formula for the Fourier coefficients of the Eisenstein series. In Section 5, we briefly recall the Weil representation for $J$. The Fourier-Jacobi expansion along $J$ of Eisenstein series associated to $P$ fixed by G(^)G($\widehat{\mathbb{Z}}$) is also explicitly computed in Section 6 and 7. Then, applying the theory of a compatible family of Eisenstein series due to Ikeda [10], [11], we give a uniform proof for the main theorems in Section 8. We also compute the $L$-functions of our lifts in Section 9, and interpret our lifts in terms of CAP representations in Section 10. Finally, Appendix is given for computing some integrals which show up in Section 7.
Acknowledgments. We would like to thank Hiroaki Narita and Jim Arthur for helpful discussions. We thank KIAS in Seoul and Waseda University in Tokyo for their incredible hospitality during this research.
2. Preliminaries
For each quadratic space $W$ or its representation matrix $A$, the corresponding symmetric pairing is denoted by $(\ast,\ast)_{W}$ or $(\ast,\ast)_{A}$.
For each integer $n\geq 3$, let $A$ be a positive definite symmetric matrix in $M_{n-2}(\mathbb{Z})$ as defined in Section 1. Put
$J_{1,n-1}=\left(\begin{array}[]{ccc}0&0&1\\ 0&-A&0\\ 1&0&0\end{array}\right),\ J_{2,n}=\left(\begin{array}[]{ccc}0&0&1\\ 0&J_{1,n-1}&0\\ 1&0&0\end{array}\right),\ J_{3,n+1}=\left(\begin{array}[]{ccc}0&0&1\\ 0&J_{2,n}&0\\ 1&0&0\end{array}\right).$
Let $V^{\prime}=\mathbb{G}^{n+2}_{a}$ be the quadratic space associated to $J_{2,n}$. Let $V=\mathbb{G}_{a}e\oplus V^{\prime}\oplus\mathbb{G}_{a}f=\mathbb{G}^{n+4}_{a}$ be the quadratic space associated to $J_{3,n+1}$ where
$e:=(1,\overbrace{0,\ldots,0}^{n+3}),\ f:=(\overbrace{0,\ldots,0}^{n+3},1)$
so that $(e,f)_{V}=\frac{1}{2}eJ_{3,n+1}{}^{t}f=\frac{1}{2}$. Let $q$ be the quadratic form on $V^{\prime}$ defined by $q(x)=\frac{1}{2}xJ_{2,n}{}^{t}x$ for $x\in V^{\prime}$. Denote by
(x,y)=q(x+y)q(x)q(y)(x,y)=q(x+y)-q(x)-q(y)
for $x,y\in V^{\prime}$ the associated bilinear form. The quadratic form $\tilde{q}$ on $V$ for $J_{3,n+1}$ satisfies $\tilde{q}(\alpha e+v^{\prime}+\beta f)=\alpha\beta+q(v^{\prime})$.
Let
$G={\rm SO}(V)={\rm SO}(3,n+1)=\{g\in{\rm SL}_{n+4}\ |\ {}^{t}gJ_{3,n+1}g=J_{3,n+1}\}$
and we also consider
${\rm SO}(V^{\prime})={\rm SO}(2,n)=\{g\in{\rm SL}_{n+2}\ |\ {}^{t}gJ_{2,n}g=J_{2,n}\}.$
The natural embedding $V^{\prime}\hookrightarrow V,\ v^{\prime}\mapsto(0,v^{\prime},0)$ yields an embedding
${\rm SO}(V^{\prime})\hookrightarrow{\rm SO}(V),\ g^{\prime}\mapsto\left(\begin{array}[]{ccc}1&0&0\\ 0&g^{\prime}&0\\ 0&0&1\end{array}\right).$
Throughout this paper, we assume $G={\rm SO}(V)$ splits at any finite place of $\mathbb{Q}$ as in Section 1.
Remark 2.1.
Our definition of $q$ is slightly different from the one in [23], which we recall in Section 4.2. Then, our ${\rm SO}(V^{\prime})$ is isomorphic over $\mathbb{R}$ to Pollack’s one but not over $\mathbb{Q}$ since any normalized orthogonal transformation matrices have to involve $\sqrt{2}$.
2.1. Siegel parabolic subgroup $P=MN$
We define
$M=\{{\rm diag}(t,m,t^{-1})\in G\ |\ t\in{\rm GL}_{1},\ m\in{\rm SO}(V^{\prime})\}\simeq{\rm GL}_{1}\times{\rm SO}(V^{\prime})$
with $\nu:M\longrightarrow{\rm GL}_{1},\ {\rm diag}(t,m,t^{-1})\mapsto t$, and
$N=\left\{n(\textbf{x}):=\left(\begin{array}[]{ccc}1&-\textbf{x}J_{2,n}&-\frac{1}{2}\textbf{x}J_{2,n}{}^{t}\textbf{x}\\ 0&1_{n+2}&{}^{t}\bf{x}\\ 0&0&1\end{array}\right)\Bigg|\ \textbf{x}\in\mathbb{G}^{n+2}_{a}\right\}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}V^{\prime},\ n(x)\mapsto x.$
Then, ${\rm diag}(t,m,t^{-1})n\in P$ acts on $x\in V^{\prime}$ by $t^{2}(m\cdot x)$ where $m\cdot x$ stands for the usual matrix multiplication.
Once a suitable Haar measure is chosen, the modulus character of $P$ is given by $\delta_{P}({\rm diag}(t,m,t^{-1}))=|t|^{n+2}$.
2.2. Another maximal parabolic subgroup and a Jacobi group
Let $V^{\prime\prime}$ be the quadratic space associated to $J_{1,n-1}$. Let
$Q=\left\{\left(\begin{array}[]{ccc}A_{1}&\ast&\ast\\ 0&m&\ast\\ 0&0&A_{2}\end{array}\right)\in G\Bigg|\ A_{1},A_{2}\in{\rm GL}_{2},\ m\in{\rm O}(V^{\prime\prime})\right\}$
be another maximal parabolic subgroup. Explicitly, if we write $Q=LU$, then
$L=\left\{\left(\begin{array}[]{ccc}\det(\gamma)^{-1}\eta\gamma\eta&0&0\\ 0&m&0\\ 0&0&\gamma\end{array}\right)\ \Bigg|\ \gamma\in{\rm GL}_{2},\ m\in{\rm SO}(V^{\prime\prime})\right\}\simeq{\rm GL}_{2}\times{\rm SO}(V^{\prime\prime})$
where $\eta=\left(\begin{array}[]{cc}-1&0\\ 0&1\end{array}\right)$ and the unipotent radical $U$ is consisting of all $v(x,y,z):=v(x,y,0)v(0,0,z)\in G$ where
$v(x,y,0)=\left(\begin{array}[]{ccccc}1&0&-yJ_{1,n-1}&-\frac{1}{2}xJ_{1,n-1}{}^{t}y&-\frac{1}{2}yJ_{1,n-1}{}^{t}y\\ 0&1&-xJ_{1,n-1}&-\frac{1}{2}xJ_{1,n-1}{}^{t}x&-\frac{1}{2}yJ_{1,n-1}{}^{t}x\\ 0&0&1_{n}&{}^{t}x&{}^{t}y\\ 0&0&0&1&0\\ 0&0&0&0&1\end{array}\right),\ v(0,0,z)=\left(\begin{array}[]{ccccc}1&0&0&-z&0\\ 0&1&0&0&z\\ 0&0&1_{n}&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{array}\right)$
for $x,y\in\mathbb{G}^{n}_{a},z\in\mathbb{G}_{a}$. Then,
$U=\{v(x,y,z)\in G\ |\ x,y\in\mathbb{G}^{n}_{a},z\in\mathbb{G}_{a}\}$
is a Heisenberg group with the center $Z=\{v(0,0,z)\ |\ z\in\mathbb{G}_{a}\}$ and we have
$v(x,y,z)v(x^{\prime},y^{\prime},z^{\prime})=v(x+x^{\prime},y+y^{\prime},z+z^{\prime}+\frac{1}{2}\langle(x,y),(x^{\prime},y^{\prime})\rangle)$
where $\langle(x,y),(x^{\prime},y^{\prime})\rangle:=xJ_{1,n-1}{}^{t}y^{\prime}-x^{\prime}J_{1,n-1}{}^{t}y$ defines a symplectic form on $XY$ with $X=\{v(x,0,0)\ |\ x\in\mathbb{G}^{n}_{a}\},\ Y=\{v(0,y,0)\ |\ y\in\mathbb{G}^{n}_{a}\}$. We identify $X$ with $V^{\prime\prime}$. Put
(2.1) $\sigma(x,y):=\langle(x,0),(0,y)\rangle=xJ_{1,n-1}{}^{t}y$
which shows up in the action of the Weil representation in our setting.
We define
$H:=\Bigg\{\ell(\gamma)\in L\ \Bigg|\ \gamma=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in{\rm GL}_{2}\Bigg\}\simeq{\rm GL}_{2},\ \ell(\gamma)\longleftrightarrow\gamma$
where
$\ell(\gamma)=\left(\begin{array}[]{ccc}\det(\gamma)^{-1}\eta\gamma\eta&0&0\\ 0&1_{n}&0\\ 0&0&\gamma\end{array}\right),\ \eta=\left(\begin{array}[]{cc}-1&0\\ 0&1\end{array}\right).$
It is easy to see that
$\ell^{-1}(\gamma)v(x,y,z)\ell(\gamma)=v(ax+cy,bx+dy,\det(\gamma)z),\ \gamma=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)\in{\rm GL}_{2},\ x,y\in\mathbb{G}^{n}_{a},\ z\in\mathbb{G}_{a}$
as in [15, (4.5),p.231]. Thus, we have the Jacobi group in our setting:
$J:=U\rtimes H^{\rm ss},\ H^{\rm ss}:=[H,H]\simeq{\rm SL}_{2}.$
3. Automorphic forms on $G={\rm SO}(3,n+1)$
Recall $G={\rm SO}(V)$. The maximal compact subgroup $K$ of $G(\mathbb{R})$ is $S({\rm O}(3)\times{\rm O}(n+1))$. We have a map $S({\rm O}(3)\times{\rm O}(n+1))\rightarrow{\rm O}(3)=({\rm SU}(2)/\mu_{2})\rtimes\{\pm 1\}$. Denote by $\mathbb{V}_{l}={\rm Sym}^{2l}(\mathbb{C}^{2})$ the $(2l+1)$-dimensional representation of $K$ that factors through ${\rm O}(3)$. We fix a basis $\{[x]^{l+v}[y]^{l-v}\}_{v=-l}^{l}$ of $V_{l}$ where $x^{k}]=\frac{x^{k}}{k!}$ and $y^{k}]=\frac{y^{k}}{k!}$ for $k\geq 0$. Modular forms on $G$ of weight $l$ are $\mathbb{V}_{l}$-valued automorphic functions $\phi$ on $G(\mathbb{A})$ that satisfy:
(1) $\phi(gk)=k^{-1}\phi(g)$ for all $g\in G(\mathbb{A})$ and $k\in K$.
(2) $\phi$ is annihilated by a special differential operator $\mathcal{D}_{l}$.
Theorem 3.1.
[23] Suppose $\phi$ is a modular form of weight $l\geq 1$ on $G$. Then
$\phi(g)=\phi_{0}(g)+\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)\geq 0,\eta\neq 0}a_{\phi}(\eta)(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),$
for $g=g_{\mathbb{f}}g_{\infty}$ in $G(\mathbb{A}_{\mathbb{f}})\times G(\mathbb{R})$, where $a_{\phi}(\eta):G(\mathbb{A}_{\mathbb{f}})\longrightarrow\mathbb{C}$ is a locally constant function. Moreover,
$\phi_{0}(m)=\Phi(m)x^{2l}+\beta(m_{f})x^{l}y^{l}+\Phi^{\prime}(m)y^{2l},$
where $\Phi$ is an automorphic function associated to a holomorphic modular form of weight $l$ on $M$, $\beta$ is a locally constant function on $M(\mathbb{A}_{\mathbb{f}})$, and $\Phi^{\prime}$ is a certain $(K\cap M)$-right translate of $\Phi$.
Here $\mathcal{W}_{\eta}(g):G(\mathbb{R})\longrightarrow\mathbb{V}_{l}$ is a generalized Whittaker function of type $\eta$ satisfying
  1. (1)
    $\mathcal{W}_{\eta}(n(x)g)=e^{i(\eta,x)}\mathcal{W}_{\eta}(g)$;
  2. (2)
    $\mathcal{W}_{\eta}(gk)=k^{-1}\mathcal{W}_{\eta}(g)$;
  3. (3)
    $D_{l}\mathcal{W}_{\eta}(g)=0$;
  4. (4)
    Suppose $\eta\in V^{\prime}(\mathbb{R})$, $\eta\neq 0$, and $(\eta,\eta)\geq 0$. For $t\in GL_{1}(\mathbb{R}),m\in{\rm SO}(V^{\prime})(\mathbb{R})$, set
    (3.1) $u_{\eta}(t,m)=it(\eta,m(iv_{1}-v_{2})).$
    We should remark that Pollack used $u_{\eta}(t,m)=it\sqrt{2}(\eta,m(iv_{1}-v_{2}))$ in [23] because of the shape of $q$ (see also Remark 2.1). Then
    (3.2) $\mathcal{W}_{\eta}(t,m)=t^{l}|t|\sum_{-l\leq v\leq l}\left(\frac{|u_{\eta}(t,m)|}{u_{\eta}(t,m)}\right)^{v}K_{v}(|u_{\eta}(t,m)|)[x^{l+v}][y^{l-v}],$
    where $K_{v}$ is the $K$-Bessel function defined by K_v(y)=12∫_0^∞e^-y(t+t^-1)/2 t^v  dtt.
Definition 3.2.
We say that $\phi$ is a cusp form of level one if it has the Fourier expansion of the form
$\phi(g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}a_{\phi}(\eta)(g_{f})\mathcal{W}_{2\pi\eta}(g_{\infty}).$
By definition, vectors $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$ are anisotropic vectors and they are called rank 2 elements in [23].
Remark 3.3.
[…]
4. Eisenstein series
In this section, we follow [23, Section 4]. For any $l>n+1$ and a Schwartz-function $\Phi_{\mathbb{f}}$ on $V(\mathbb{A}_{\bf f})$, we define a section $f$ of ${\rm Ind}^{G(\mathbb{A})}_{P(\mathbb{A})}\,|\nu|^{s-\frac{n+2}{2}}$ (normalized induction) by
$f_{l}(g,\Phi_{\mathbb{f}},s)=f_{{\rm fte}}(g_{\mathbb{f}},\Phi_{\mathbb{f}},s)f_{l,\infty}(g_{\infty},s),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})$
where
$f_{{\rm fte}}(g_{\mathbb{f}},\Phi_{\mathbb{f}},s)=\int_{{\rm GL}_{1}(\mathbb{A}_{\mathbb{f}})}|t|^{s}\Phi_{\mathbb{f}}(tg_{\mathbb{f}}e)dt,\ g_{\mathbb{f}}\in G(\mathbb{A}_{\mathbb{f}})$
and $f_{l,\infty}(g_{\infty},s)$ is the $V_{l}={\rm Sym}^{2l}(\mathbb{C}^{2})$-valued, $K_{\infty}$-equivalent section of ${\rm Ind}^{G(\mathbb{R})}_{P(\mathbb{R})}\,|\nu|^{s-\frac{n+2}{2}}$. See Section 4.2 for a precise definition. Then, we define the associated Eisenstein series as
[…]
By definition, $E_{l}(g,\Phi_{\mathbb{f}},s)=E_{l}(g_{\infty},g_{\mathbb{f}}\cdot\Phi_{\mathbb{f}},s)$ where $g_{\mathbb{f}}\cdot\Phi_{\mathbb{f}}$ is defined by $g_{\mathbb{f}}\cdot\Phi_{\mathbb{f}}(v)=\Phi_{\mathbb{f}}(g^{-1}_{\mathbb{f}}v)$ for $v\in V(\mathbb{A})$. It converges absolutely for $Re(s)>n+2$. When $s=l+1$ and $l+1>n+2$, $E_{l}(g,\Phi_{\mathbb{f}},s=l+1)$ gives rise to a modular form of weight $l$, and it has the Fourier expansion as
(4.1) $E_{l}(g,\Phi_{\mathbb{f}})=E_{0}(g)+\sum_{0\neq\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)\geq 0}a_{E_{l}}(\eta)(g_{\mathbb{f}}\cdot\Phi_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty})$
[…]
(4.2) $E_{l}(g,\Phi_{\mathbb{f}})=E_{0}(g)+\sum_{0\neq\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)\geq 0}a_{E_{l}}(\eta)(\Phi_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty})$
for $g=\gamma k_{\mathbb{f}}g_{\infty}\in G(\mathbb{A})=G(\mathbb{Q})(G(\widehat{\mathbb{Z}})\times G(\mathbb{R}))$, since $k_{\mathbb{f}}\cdot\Phi_{\mathbb{f}}=\Phi_{\mathbb{f}}$.
We compute explicitly $a_{E_{l}}(\eta):=a_{E_{l}}(\eta)(\Phi_{\mathbb{f}})$. If $\eta$ is of rank 2, i.e., $q(\eta)\neq 0$,
(4.3) $a_{E_{l}}(\eta)(\Phi_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty})=J(s,\eta,f_{l})=\int_{V^{\prime}(\mathbb{A})}\psi((\eta,x))f_{l}(wn(x),\Phi_{\mathbb{f}},s)\,dx.$
where $\psi=\otimes^{\prime}_{p}\psi_{p}:\mathbb{Q}\backslash\mathbb{A}\longrightarrow\mathbb{C}^{\times}$ is the standard additive character. Write
$J(s,\eta,f_{l})=\prod_{p}J(s,\eta,\Phi_{p}),\ J(s,\eta,\Phi_{p}):=\int_{V^{\prime}(\mathbb{Q}_{p})}\psi_{p}((\eta,x))f_{l}(wn(x),\Phi_{\mathbb{f}},s)\,dx$
where $w\in G={\rm SO}(V)$ is defined by $w\cdot e=f$, $w\cdot f=e$, and $w|_{V^{\prime}}={\rm id}_{V^{\prime}}$. Notice that $J(s,\eta,\Phi_{p})$ is also defined for $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$. Further, by [23, p.633, line -6], $J(s,\eta,\Phi_{p})=\displaystyle\int_{{\rm GL}_{1}(\mathbb{Q}_{p})}\int_{V^{\prime}(\mathbb{Q}_{p})}\psi_{p}((\eta,x))\Phi_{p}(t,tx,-tq^{\prime}(x)))|t|^{s}_{p}\,dxdt$. Since $t$ runs over $\mathbb{Z}^{\times}_{p}$, $\Phi_{p}(t,tx,-tq^{\prime}(x)))$ is preserved under the transformation $x\mapsto x+x_{0}$ for any $x_{0}\in V^{\prime}(\mathbb{Z}_{p})$. By the change of variables, $J(s,\eta,\Phi_{p})=\psi((\eta,x_{0}))J(s,\eta,\Phi_{p})$ for any $x_{0}\in V^{\prime}(\mathbb{Z}_{p})$. Thus, $J(s,\eta,\Phi_{p})=0$ if $\eta\not\in V^{\prime}(\mathbb{Z}_{p})$.
The following lemma is an analogue of [16, p.591, Lemma 3.4] and it is proved by the same argument therein:
Lemma 4.1.
Keep the notation as above. For each $g_{p}={\rm diag}(t_{p},m_{p},t_{p}^{-1})n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$, it holds
$J(s,\eta,g_{p}\cdot\Phi_{p})=|t_{p}|^{\frac{n+2}{2}}_{p}J(s,g_{p}^{-1}\eta,\Phi_{p}).$
4.1. $p$-adic computation: Siegel series
Let $\psi:\mathbb{Q}^{\times}_{p}\longrightarrow\mathbb{C}^{\times}$ be the standard additive character and henceforth we simply write it by $\psi(x)=e^{2\pi ix}$ (though $x$ should be the fractional part of $x$ to be more precise). For $\Phi_{p}={\rm char}_{V^{\prime}(\mathbb{Z}_{p})}$ and $\eta\in V^{\prime}(\mathbb{Q}_{p})$ with $q(\eta)\neq 0$, we will compute the Siegel series:
$J(s,\eta,\Phi_{p})=\sum_{r=0}^{\infty}p^{-rs}\int_{p^{-r}V^{\prime}(\mathbb{Z}_{p})}\psi((\eta,x))\text{char}(p^{r}q(x)\in\mathbb{Z}_{p})\,dx.$
We denote the integral by $B_{r,\eta}$. Then by the change of variables,
$B_{r,\eta}=\int_{p^{-r}V^{\prime}(\mathbb{Z}_{p})}\psi((\eta,x))\text{char}(p^{r}q(x)\in\mathbb{Z})\,dx=p^{r(n+2)}\int_{V^{\prime}(\mathbb{Z}_{p})\atop q(x)\equiv 0\,\text{mod $p^{r}$}}\psi\Big(\frac{(\eta,x)}{p^{r}}\Big)\,dx.$
Set $x=u+y$, where $u\in V^{\prime}(\mathbb{Z}_{p}/p^{r}\mathbb{Z}_{p})$, and $y\in V^{\prime}(p^{r}\mathbb{Z}_{p})$. If $y\equiv 0$ (mod $p^{r}$), $q(y)\equiv 0$ (mod $p^{r}$). It follows from this that, upon making the change of variables in $y$, the factor $p^{-r(n+2)}$ arises, cancelling the factor $p^{r(n+2)}$ obtained earlier. Thus
(4.4) $B_{r,\eta}=\sum_{u\in V^{\prime}(\mathbb{Z}/p^{r}\mathbb{Z})\atop q(u)\equiv 0\,\text{(mod $p^{r}$)}}\psi\Big(\frac{(\eta,u)}{p^{r}}\Big).$
When $r=0$, we have the integral
$\int_{V^{\prime}(\mathbb{Z}_{p})}\psi((\eta,x))\,dx.$
Then it is $\begin{cases}1,&\text{if $\eta\in V^{\prime}(\mathbb{Z}_{p})$}\\ 0,&\text{otherwise}\end{cases}$.
If $r\geq 1$, we need to compute
$\sum_{u\in V^{\prime}(\mathbb{Z}/p^{r}\mathbb{Z})\atop q(u)\equiv 0\,\text{(mod $p^{r}$)}}\psi\Big(\frac{(\eta,u)}{p}\Big).$
We need to divide into three cases:
Theorem 4.2.
Suppose $n$ is even. Let $p^{k}\|\eta$, and $p^{2k+k^{\prime}}\|q(\eta)$. Then
$J(s,\eta,\Phi_{p})=(1-p^{m-1-s})Q_{\eta,p}(p^{m-s}),$
where $Q_{\eta,p}(X)\in\mathbb{Z}[X]$ is a polynomial of degree $2k+k^{\prime}$, and it satisfies the functional equation
$X^{2k+k^{\prime}}Q_{\eta,p}(X^{-1})=Q_{\eta,p}(X).$
Let
$\widetilde{Q}_{\eta,p}(X)=X^{2k+k^{\prime}}Q_{\eta,p}(X^{-2}).$
Then $\widetilde{Q}_{\eta,p}(X^{-1})=\widetilde{Q}_{\eta,p}(X)$.
Special case: $k=0,k^{\prime}=0$. Then $J(s,\eta,\Phi_{p})=1-p^{m-1-s}$. If $k=0$, $k^{\prime}=1$, $J(s,\eta,\Phi_{p})=1+(p^{m}-p^{m-1})p^{-s}-p^{2m-1-2s}=(1-p^{m-1-s})(1+p^{m-s})$. Hence $Q_{\eta,p}(p^{m-s})=1+p^{m-s}$. More generally, if $k=0$, $Q_{\eta,p}(p^{m-s})=\sum_{a=0}^{k^{\prime}}(p^{m-s})^{a}$.
Theorem 4.3.
Let $n$ be odd and $p>2$. Let $p^{k}\|\eta$ and $p^{2k+k^{\prime}}\|q(\eta)$, and $q(\eta)_{1}=\frac{q(\eta)}{p^{2k+k^{\prime}}}$. Let $\chi_{\eta}(p)=\begin{cases}(\frac{q(\eta)_{1}}{p}),&\text{if $2k+k^{\prime}$ even}\\ 0,&\text{if $2k+k^{\prime}$ odd}\end{cases}$. Then
$J(s,\eta,\Phi_{p})=\frac{1-p^{2(m-s)}}{1-\chi_{\eta}(p)p^{m-s}}Q_{\eta,p}(p^{m-s}),$
where $Q_{\eta,p}(X)\in\mathbb{Z}[X]$ is a polynomial of degree $2\lfloor\frac{2k+k^{\prime}}{2}\rfloor$, and it satisfies the functional equation
$p^{l}X^{2l}Q_{\eta,p}(p^{-1}X^{-1})=Q_{\eta,p}(X),$
where $2k+k^{\prime}=2l$ or $2l+1$.
Let $\widetilde{Q}_{\eta,p}(X)=X^{-l}Q_{\eta,p}(p^{-\frac{1}{2}}X)$. Then
$\widetilde{Q}_{\eta,p}(X^{-1})=\widetilde{Q}_{\eta,p}(X).$
Special case: $k=0,k^{\prime}=0$. Then $J(s,\eta,\Phi_{p})=1+\Big(\frac{q(\eta)}{p}\Big)p^{m-s}$. If $k=0$, $k^{\prime}=1$, $J(s,\eta,\Phi_{p})=1-p^{2(m-s)}$, and so $Q_{\eta,p}(p^{m-s})=1$. If $k=0$, $k^{\prime}=2$, $Q_{\eta,p}(p^{m-s})=1-\Big(\frac{q(\eta)_{1}}{p}\Big)p^{m-s}+p^{2(m-s)+1}$.
More generally, if $k=0$, $k^{\prime}=2l$ even, (let $\varepsilon=\Big(\frac{q(\eta)_{1}}{p}\Big)$), Then
$J(s,\eta,\Phi_{p})=1+\sum_{r=1}^{l}p^{2r(m-s)}(p^{r}-p^{r-1})+\varepsilon p^{(2l+1)(m-s)+l}=(1+\varepsilon p^{m-s})Q_{\eta,p}(p^{m-s}),$
where $Q_{\eta,p}(p^{m-s})=\sum_{i=0}^{l}p^{2i(m-s)+i}-\varepsilon\sum_{i=1}^{l}p^{(2i-1)(m-s)+i-1}.$
If $k=0$, $k^{\prime}=2l+1$ odd ($l\geq 1$),
$J(s,\eta,\Phi_{p})=1+\sum_{r=1}^{l}p^{2r(m-s)}(p^{r}-p^{r-1})-p^{(2l+2)(m-s)+l}=(1-p^{2(m-s)})Q_{\eta,p}(p^{m-s}),$
where $Q_{\eta,p}(p^{m-s})=1+\sum_{i=1}^{l}p^{2i(m-s)+i}.$
Recall that $(\tfrac{n}{2})=0$ if $2|n$ or $n\equiv 3$ (mod 4). For $n\equiv 1$ (mod 4), $(\tfrac{n}{2})=\begin{cases}1,&\text{if $n\equiv 1$ (mod 8)}\\ -1,&\text{if $n\equiv 5$ (mod 8)}\end{cases}$.
Theorem 4.4.
Let $n$ be odd and $p=2$. Let $2^{k}\|\eta$ and $2^{2k+k^{\prime}}\|q(\eta)$. Let $l=\lfloor\frac{2k+k^{\prime}}{2}\rfloor$. Let $q(\eta)_{1}=q(\eta)2^{-2k-k^{\prime}}$. Let $\chi_{\eta}(2)=\begin{cases}(\frac{q(\eta)_{1}}{2}),&\text{if $2k+k^{\prime}$ even}\\ 0,&\text{if $2k+k^{\prime}$ odd}\end{cases}$. Then
$J(s,\eta,\Phi_{2})=\frac{1-2^{2(m-s)}}{1-\chi_{\eta}(2)2^{m-s}}Q_{\eta,p}(2^{m-s}),$
where $Q_{\eta,p}(X)\in\mathbb{Z}[X]$ is a polynomial of degree $\alpha$, and $\alpha=\begin{cases}2l,&\text{if $\chi_{\eta}(2)=0$}\\ 2l+2,&\text{if $\chi_{\eta}(2)=\pm 1$}\end{cases}$. And $Q_{\eta,p}(X)$ satisfies the functional equation
$2^{\frac{\alpha}{2}}X^{\alpha}Q_{\eta,p}(2^{-1}X^{-1})=Q_{\eta,p}(X).$
Since the proofs are very long, we will postpone the proofs until Section 11.
4.2. Archimedean computation
For the archimedean place, recall some definitions. The $K$-Bessel function $K_{v}(y)$ is defined as
$K_{v}(y)=\frac{1}{2}\int_{0}^{\infty}e^{-y(t+t^{-1})/2}t^{v}\frac{dt}{t}.$
It satisfies $K_{-v}(y)=K_{v}(y)$. The maximal compact subgroup $K$ of ${\rm SO}(V)(\mathbb{R})$ is $S({\rm O}(3)\times{\rm O}(n+1))$. Denote by $\mathbb{V}_{l}=Sym^{2l}(\mathbb{C}^{2})$, the $(2l+1)$-dimensional representation of $K$ that factors through ${\rm O}(3)$, and let $\{X^{2l},X^{2l-1}Y,...,Y^{2l}\}$ be a basis of $\mathbb{V}_{l}$.
Recall the notations in [23]: $V^{\prime}=V_{2}\oplus V_{n}$ with the quadratic form $q(x,y)=q_{2}(x)-q_{n}(y)$, where $(V_{2},q_{2})$ and $(V_{n},q_{n})$ are 2-dimensional ($n$-dimensional, resp.) quadratic space with positive definite quadratic forms. Let $\{v_{1},v_{2}\}$ be an orthonormal basis of $V_{2}(\mathbb{R})$. Note that this quadratic form is different from the one in Section 2. It accounts for the appearance of $\sqrt{2}$.
Let $f_{l}(g,s)$ be the $\mathbb{V}_{l}$-valued section. For $x\in V^{\prime}(\mathbb{R})$, let $x=(x_{2},x_{n})$, where $x_{2}\in V_{2},x_{n}\in V_{n}$. Then
[…]
where $p_{V_{3}}:V=V_{3}\oplus V_{n+1}\longrightarrow V_{3}$ is the orthogonal projection. Then
[…]
and
[…]
Consider, for $\eta\in V^{\prime}(\mathbb{R})$ with $q(\eta)>0$,
$J(s,f_{l})=\int_{V^{\prime}(\mathbb{R})}e^{-2\pi i(\eta,x)}f_{l}(wn(x),s)\,dx.$
[…]
Lemma 4.5.
[23] If $b>0$,
$\left(aX^{2}+2bXY-\bar{a}Y^{2}\right)^{l}=\sum_{v=0}^{l}\begin{pmatrix}l\\ v\end{pmatrix}2^{l-v}\delta_{v0}^{\frac{1}{2}}(XY)^{l-v}(|a|^{2}+b^{2})^{\frac{l-v}{2}}$
$\phantom{xxxxxxxxxxxxx}\cdot{}_{2}F_{1}\left(\frac{v-l}{2},\frac{v+l+1}{2};v+1,\frac{|a|^{2}}{|a|^{2}+b^{2}}\right)\left(a^{v}X^{2v}+(-\bar{a})^{v}Y^{2v}\right),$
where $\delta_{v0}^{\frac{1}{2}}=\begin{cases}\frac{1}{2},&\text{if $v=0$}\\ 1,&\text{otherwise}\end{cases}$.
Here we use the identity
$_{2}F_{1}(a^{\prime},b^{\prime};c^{\prime},z)=(1-z)^{-a^{\prime}}{}_{2}F_{1}(a^{\prime},c^{\prime}-b^{\prime};c^{\prime},\frac{z}{z-1}).$
So
$(|a|^{2}+b^{2})^{\frac{l-v}{2}}{}_{2}F_{1}\left(\frac{v-l}{2},\frac{v+l+1}{2};v+1,\frac{|a|^{2}}{|a|^{2}+b^{2}}\right)=b^{l-v}{}_{2}F_{1}\left(\frac{v-l}{2},\frac{v-l}{2}+\frac{1}{2};v+1,-\frac{|a|^{2}}{b^{2}}\right).$
So
$J(s,f_{l})=(2\sqrt{2})^{-l}\int_{V^{\prime}(\mathbb{R})}e^{-2\pi i(\eta,x)}\tau(\sqrt{2}x_{n},\sqrt{2}x_{2})^{-s-l}\left(aX^{2}+2bXY-\bar{a}Y^{2}\right)^{l}\,dx.$
[…]
$\Phi_{v}(s)=\int_{V^{\prime}(\mathbb{R})}e^{-2\pi i(\eta,x)}\tau(\sqrt{2}x_{n},\sqrt{2}x_{2})^{-s-v}a^{v}{}_{2}F_{1}\left(\frac{v-l}{2},\frac{l+v+1}{2};v+1,\frac{|a|^{2}}{|a|^{2}+b^{2}}\right)\,dx$
$\phantom{xxxx}=\int_{V^{\prime}(\mathbb{R})}e^{-2\pi i(\eta,x)}(|a|^{2}+b^{2})^{-\frac{s+l}{2}}a^{v}b^{l-v}{}_{2}F_{1}\left(\frac{v-l}{2},\frac{v-l}{2}+\frac{1}{2};v+1,-\frac{|a|^{2}}{b^{2}}\right)\,dx.$
Let
$\widetilde{\Phi}_{v}(s)=\int_{V^{\prime}(\mathbb{R})}e^{-2\pi i(\eta,x)}\tau(\sqrt{2}x_{n},\sqrt{2}x_{2})^{-s-v}(-\bar{a})^{v}{}_{2}F_{1}\left(\frac{v-l}{2},\frac{l+v+1}{2};v+1,\frac{|a|^{2}}{|a|^{2}+b^{2}}\right)\,dx\\$
Then $\Phi_{0}(s)=\widetilde{\Phi}_{0}(s)$, and
$J(s,f_{l})=(2\sqrt{2})^{-l}\sum_{v=1}^{l}\begin{pmatrix}l\\ v\end{pmatrix}2^{l-v}\Phi_{v}(s)X^{l+v}Y^{l-v}+(2\sqrt{2})^{-l}\sum_{v=1}^{l}\begin{pmatrix}l\\ v\end{pmatrix}2^{l-v}\tilde{\Phi}_{v}(s)X^{l-v}Y^{l+v}$
$\phantom{xxxxxxx}+(2\sqrt{2})^{-l}2^{l}\Phi_{0}(s)X^{l}Y^{l}.$
Pollack [23] proved the following when $s=l+1$,
Theorem 4.6.
Let $v\geq 0$.
$\Phi_{v}(l+1)=\frac{(2\pi)^{-\frac{n+2}{2}}2^{-(l-v-1-\frac{n}{2})}\Gamma(v+1)}{\Gamma(l+v+1)\Gamma(l-\frac{n}{2}+1)}q(\eta)^{l-\frac{n}{2}}\left(\frac{|(2\pi\eta,v_{1}+iv_{2})|}{(2\pi\eta,v_{1}+iv_{2})}\right)^{v}K_{v}(\sqrt{2}|(2\pi\eta,v_{1}+iv_{2})|).$
We see that since $K_{-v}(x)=K_{v}(x)$,
$\widetilde{\Phi}_{v}(l+1)=\frac{(2\pi)^{-\frac{n+2}{2}}2^{-(l-v-1-\frac{n}{2})}\Gamma(v+1)}{\Gamma(l+v+1)\Gamma(l-\frac{n}{2}+1)}q(\eta)^{l-\frac{n}{2}}\left(\frac{|(2\pi\eta,v_{1}+iv_{2})|}{\overline{(2\pi\eta,v_{1}+iv_{2})}}\right)^{v}K_{v}(\sqrt{2}|(2\pi\eta,v_{1}+iv_{2})|).$
Hence
Theorem 4.7.
[23]
$J(l+1,f_{l})=\pi^{2l+1-\frac{n}{2}}q(\eta)^{l-\frac{n}{2}}\mathcal{W}_{2\pi\eta}(1),$
where for $\eta\in V^{\prime}(\mathbb{R})$, and $t\in GL_{1}(\mathbb{R}),m\in SO(V^{\prime})(\mathbb{R})$,
$u_{\eta}(t,m)=\sqrt{2}ti(\eta,m(iv_{1}-v_{2})),$
$\mathcal{W}_{2\pi\eta}(t,m)=t^{l}|t|\sum_{v=-l}^{l}\left(\frac{|u_{\eta}(t,m)|}{u_{\eta}(t,m)}\right)^{v}K_{v}(|u_{\eta}(t,m)|)X^{l-v}Y^{l+v}.$
We can write
$\mathcal{W}_{2\pi\eta}(t,m)=t^{l}|t|\sum_{v=1}^{l}\left(\frac{|u_{\eta}(t,m)|}{u_{\eta}(t,m)}\right)^{v}K_{v}(|u_{\eta}(t,m)|)X^{l-v}Y^{l+v}$
$\phantom{xxxxx}+t^{l}|t|\sum_{v=1}^{l}\left(\frac{|u_{\eta}(t,m)|}{\overline{u_{\eta}(t,m)}}\right)^{v}K_{v}(|u_{\eta}(t,m)|)X^{l+v}Y^{l-v}+t^{l}|t|K_{0}(|u_{\eta}(t,m)|)X^{l}Y^{l}.$
After combining $p$-adic and archimedean computations, we have
Theorem 4.8 ($n$ even).
(The rank 2 case) For $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$, we have
$a_{E}(\eta)=C_{l,n}\prod_{p}Q_{\eta,p}(p^{l-\frac{n}{2}})=C_{l,n}q(\eta)^{\frac{l-\frac{n}{2}}{2}}\prod_{p}\widetilde{Q}_{\eta,p}(p^{\frac{l-\frac{n}{2}}{2}})$
where $C_{l,n}=\displaystyle\frac{\pi^{2l+1-\frac{n}{2}}}{\zeta(l+1-\frac{n}{2})}$.
One corollary of our explicit computation of the Siegel series is that $a_{E}(\eta)$ have the bounded denominator since $Q_{\eta,p}(X)\in\mathbb{Z}[X]$.
For $n$ odd, we need some modification. Write $q(\eta)=\mathfrak{d}_{\eta}\mathfrak{f}_{\eta}^{2}$ such that $\mathfrak{d}_{\eta}$ is the absolute discriminant of $\mathbb{Q}\left(\sqrt{\epsilon q(\eta)}\right)/\mathbb{Q}$ as in the introduction. Let $\chi_{\eta}$ be the primitive Dirichlet character corresponding to $\mathbb{Q}\left(\sqrt{\epsilon q(\eta)}\right)/\mathbb{Q}$. We also need the following the functional equation of Dirichlet $L$-function:
$L(l-\tfrac{n-1}{2},\chi_{\eta})=(-1)^{l+n}(2\pi)^{l-\frac{n+1}{2}}\pi\frac{\mathfrak{d}_{\eta}^{\frac{n}{2}-l}}{(l-\frac{n+1}{2})!}L(\tfrac{n+1}{2}-l,\chi_{\eta}).$
Theorem 4.9 ($n$ odd).
For rank 2 case, namely, $q(\eta)>0$, we have
$a_{E}(\eta)=\frac{L(l-\frac{n-1}{2},\chi_{\eta})}{\zeta(2l-n+1)}\pi^{2l+1-\frac{n}{2}}q(\eta)^{l-\frac{n}{2}}\prod_{p}Q_{\eta,p}(p^{\frac{n-1}{2}-l})$
$\phantom{xxxx}=C_{l,n}^{\prime}L(\tfrac{n+1}{2}-l,\chi_{\eta})\mathfrak{f}_{\eta}^{l-\frac{n}{2}}\prod_{p}\widetilde{Q}_{\eta,p}(p^{l-\frac{n}{2}}),$
where $C_{l,n}^{\prime}=\frac{(-1)^{\frac{3n+1}{2}}2^{l-\frac{n-1}{2}}\pi^{l+\frac{1}{2}}(2l-n+1)!}{B_{2l-n+1}(l-\frac{n+1}{2})!}$ and $B_{2l-n+1}$ is the Bernoulli number.
We can write $a_{E}(\eta)=(*)\pi^{l+\frac{1}{2}}L(\frac{n-1}{2}-l,\chi_{\eta})\prod_{p}Q_{\eta,p}(p^{l-\frac{n+1}{2}})$, where ()(*) is a rational number. Since $L(\frac{n-1}{2}-l,\chi_{\eta})\in\mathbb{Q}$ and $Q_{\eta,p}(X)\in\mathbb{Z}[X]$, $a_{E}(\eta)$ have the bounded denominator.
5. The Weil representation of the Jacobi group $J$
We refer to [9] for the basics of the Weil representation. Let $\psi:\mathbb{Q}\backslash\mathbb{A}\longrightarrow\mathbb{C}^{\times}$ be the standard additive character (cf [15, Section 4.2]). For each $S\in\mathbb{Q}^{\times}$, let $\psi_{S}=\psi(S\ast)=\otimes^{\prime}_{p}\psi_{S,p}$ where we also regard it with an additive character on $Z(\mathbb{A})$ via the identity $Z(\mathbb{A})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{A},\ v(0,0,z)\mapsto z$.
Let $\widetilde{H^{\rm ss}(\mathbb{A})}$ be the metaplectic covering of $H^{\rm ss}(\mathbb{A})$. Let $\omega_{\psi_{S}}:=\otimes^{\prime}\omega_{S,p}=\omega_{S,\mathbb{f}}\otimes\omega_{S,\infty}:\widetilde{J(\mathbb{A})}:=\widetilde{H^{\rm ss}(\mathbb{A})}\ltimes U(\mathbb{A})\longrightarrow{\rm Aut}_{\mathbb{C}}(\mathcal{S}(X(\mathbb{A})))$ be the Weil representation associated to $\psi_{S}$ acting on the Schwartz space $\mathcal{S}(X(\mathbb{A}))$ for $X(\mathbb{A})$. Explicitly, for each place $p\leq\infty$ and $\Phi=\otimes^{\prime}_{p\leq\infty}\Phi_{p}\in\mathcal{S}(X(\mathbb{A}))$, it is given by, for $\varepsilon\in\{\pm 1\}$, $v(x,y,z)\in U(\mathbb{Q}_{p}),\ t\in X(\mathbb{Q}_{p}),\ a\in\mathbb{Q}^{\times}_{p}$, and $b\in\mathbb{Q}_{p}$,
  1. (1)
    $\omega_{S,p}(v(x,y,z))\Phi_{p}(t)=\Phi_{p}(t+x)\psi_{S,p}\Big(z+\sigma(t,y)+\frac{1}{2}\sigma(x,y)\Big).$
  2. (2)
    $\omega_{S,p}((\ell(\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}),\varepsilon))\Phi_{p}(t)=\varepsilon^{n}\displaystyle\frac{\gamma_{p}(1)}{\gamma_{p}(S^{n})}|a|^{\frac{n}{2}}\Phi_{p}(at)$,
  3. (3)
    $\omega_{S,p}(\ell(\begin{pmatrix}1&b\\ 0&1\end{pmatrix}),\varepsilon))\Phi_{p}(t)=\varepsilon^{n}\psi_{S,p}(\sigma(t,t)b)\Phi_{p}(t),\ b\in\mathbb{Q}_{p},\ t\in X(\mathbb{Q}_{p})$.
  4. (4)
    $\omega_{S,p}(\ell(\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right)),\varepsilon))\Phi_{p}(t)=\varepsilon^{n}\gamma_{p}(1)^{-n}(F_{S}\Phi_{p})(-t),\ t\in X(\mathbb{Q}_{p})$, (F_SΦ_p)(t)=∫X(Qp) Φp(x)ψS,p(σ(t,x))dx
where $\gamma_{p}(a)$ stands for the Weil constant associated to $\psi_{S,p}$and $dx$ means the Haar measure on $X(\mathbb{Q}_{p})$ which is self-dual with respect to the Fourier transform $F_{S}$.
Recall that $X(\mathbb{R})=V^{\prime\prime}(\mathbb{R})$ is a quadratic space with signature $(p,q)=(1,n)$. By direct computation, it is easy to see that the $\widetilde{{\rm SO}(2)}-$type of $\omega_{S,\infty}$ (for $S\in\mathbb{R}^{\times}$) is given by
(5.1) ${\rm sgn}(S)\Big(\frac{p-q}{2}+m\Big)={\rm sgn}(S)\Big(\frac{1-n}{2}+m\Big),\ m\in\mathbb{Z}_{\geq 0}.$
For each $\Phi\in\mathcal{S}(X(\mathbb{A}))$, we define the theta function
$\Theta_{\psi_{S}}(v(x,y,z)\widetilde{h};\Phi)$ $:=$ $\sum_{\xi\in X(\mathbb{Q})}\omega_{\psi_{S}}(v(x,y,z)\widetilde{h})\Phi(\xi),\ v(x,y,z)\in U(\mathbb{A}),\ \widetilde{h}\in\widetilde{H^{\rm ss}(\mathbb{A})}$
$=$ $\sum_{\xi\in X(\mathbb{Q})}(\omega_{\psi_{S}}(\widetilde{h})\Phi)(x+\xi)\psi(S\sigma(\xi,y))\psi\Big(S(z+\frac{1}{2}\sigma(x,y))\Big).$
6. Fourier-Jacobi expansions of Eisenstein series along $J$: the general case
In this section we will prove that any Fourier Jacobi coefficient of Eisenstein series are Eisenstein series. We refer to [9] for basics of this topic.
6.1. Double coset decomposition
Since $n\geq 3$, the relative root system of $G$ over $\mathbb{Q}$ is of type $B_{3}$, and we denote the positive roots by $\{e_{1}\pm e_{2},e_{1}\pm e_{3},e_{2}\pm e_{3},e_{1},e_{2},e_{3}\}$, and let $\Delta=\{e_{1}-e_{2},e_{2}-e_{3},e_{3}\}$ be the set of simple roots, where $e_{i}\pm e_{j}$ has multiplicity one, and $e_{i}$ has multiplicity $8m$.
[…]
In terms of group elements, we can take coset representatives as
$\xi_{1}=\left(\begin{array}[]{ccccccc}0&0&1&0&0&0&0\\ 1&0&0&0&0&0&0\\ 0&1&0&0&0&0&0\\ 0&0&0&I_{n-2}&0&0&0\\ 0&0&0&0&0&1&0\\ 0&0&0&0&0&0&1\\ 0&0&0&0&1&0&0\end{array}\right),\quad\xi_{0}=\left(\begin{array}[]{ccccccc}0&0&0&0&0&1&0\\ 1&0&0&0&0&0&0\\ 0&0&1&0&0&0&0\\ 0&0&0&I_{n-2}&0&0&0\\ 0&0&0&0&1&0&0\\ 0&0&0&0&0&0&1\\ 0&1&0&0&0&0&0\end{array}\right).$
6.2. Fourier-Jacobi expansions of Eisenstein series
We identify $H^{{\rm ss}}$ with $SL_{2}$ via the map $\ell$. Let $B$ is the Borel subgroup of $SL_{2}$ which consists of upper-triangular matrices. Let $\omega$ be a unitary character of $\mathbb{Q}^{\times}\backslash\mathbb{A}^{\times}$ and $s\in\mathbb{C}$. Let $\mathbb{K}={\rm SL}_{2}(\widehat{\mathbb{Z}})\times{\rm SO}(2)$ be the standard maximal compact subgroup of ${\rm SL}_{2}(\mathbb{A})$. Let $\widetilde{\mathbb{K}}$ be the pullback of $\mathbb{K}$ to $\widetilde{{\rm SL}_{2}(\mathbb{A})}$ under the surjection $\widetilde{{\rm SL}_{2}(\mathbb{A})}\longrightarrow{\rm SL}_{2}(\mathbb{A})$.
We denote by $I(\omega,s)$, the degenerate principal series representation of $G(\mathbb{A})$ consisting of any function $f:G(\mathbb{A})\longrightarrow\mathbb{C}$ such that
$f(pg)=\delta^{\frac{1}{2}}_{P}(p)|\nu(p)|^{s}_{\mathbb{A}}\omega(\nu(p))f(g)$
for any $p\in P(\mathbb{A})$ and any $g\in G(\mathbb{A})$. We also define the space $I_{1}(\omega,s)$ consisting of any smooth, $\mathbb{K}$-finite function $f:{\rm SL}_{2}(\mathbb{A})\longrightarrow\mathbb{C}$ such that
$f(pg)=\delta^{\frac{1}{2}}_{B}(p)|a|^{s}_{\mathbb{A}}\omega(a)f(g)$
for any $p=\left(\begin{array}[]{cc}a&b\\ 0&a^{-1}\end{array}\right)\in B(\mathbb{A})$ and any $g\in{\rm SL}_{2}(\mathbb{A})$. Here $\delta^{\frac{1}{2}}_{B}(p)=|a|_{\mathbb{A}}$ for $p=\left(\begin{array}[]{cc}a&b\\ 0&a^{-1}\end{array}\right)\in B(\mathbb{A})$. Similarly we also define the space $\widetilde{I}_{1}(\omega,s)$ consisting of any smooth, $\widetilde{\mathbb{K}}-$finite function $f:\widetilde{{\rm SL}_{2}(\mathbb{A})}\longrightarrow\mathbb{C}$ such that
(6.1) $f(\widetilde{p}g)=\varepsilon\frac{\gamma(1)}{\gamma(a)}\delta^{\frac{1}{2}}_{B}(p)|a|^{s}_{\mathbb{A}}\omega(a)f(g)$
for any $\widetilde{p}=(p,\varepsilon)\in\widetilde{B(\mathbb{A})}$ with $p=\left(\begin{array}[]{cc}a&b\\ 0&a^{-1}\end{array}\right)$ and any $g\in\widetilde{{\rm SL}_{2}(\mathbb{A})}$. Here $\gamma:=\prod_{p}\gamma_{p}$ is the global Weil constant and $\varepsilon$ in the RHS of (6.1) is understood as multiplying +1+1 if $\varepsilon\in\{(t_{v})_{v}\in\bigoplus_{v}\{\pm 1\}\ |\ \prod_{v}t_{v}=1\}$, 1-1 otherwise.
For any section $f\in I(\omega,s)$, we define the Eisenstein series on $G(\mathbb{A})$ of type $(\omega,s)$ by
[…]
For each $S\in\mathbb{Q}^{\times}$, let
[…]
For each $S\in\mathbb{Q}^{\times}$ and $\Phi\in\mathcal{S}(X(\mathbb{A}))$, put
[…]
When ${\rm dim}\hskip 0.56905ptX=n$ is even, $\widetilde{H^{{\rm ss}}(\mathbb{A})}$ is canonically isomorphic to ${\rm SL}_{2}(\mathbb{A})\times\{\pm 1\}$ and thus, we can define
[…]
Let $\iota_{{\rm SL}_{2}}=\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right)$ be the Weyl element of ${\rm SL}_{2}$. The following theorem is proved by a standard way (see the proof of [9, p.627, Theorem 3.2]) with the double coset decomposition in Section 6.1. Thus, we omit the proof.
Theorem 6.1.
Keep the notation above. Assume that $\Phi\in\mathcal{S}(X(\mathbb{A}))$ is $\widetilde{\mathbb{K}}-$finite. For ${\rm Re}(s)\gg 0$,
  1. (1)
    when ${\rm dim}\hskip 0.56905ptX=n$ is even, $R(h;f,\Phi):=\displaystyle\int_{\mathbb{V}(\mathbb{A})}f(\xi_{0}\cdot v\cdot\iota_{{\rm SL}_{2}}\cdot h)\overline{\omega_{S}(v(\iota_{SL_{2}}\cdot h))\Phi(0)}dv$ is a section of $I_{1}(\omega,s)$,
  2. (2)
    when ${\rm dim}\hskip 0.56905ptX=n$ is odd, $R(\widetilde{h};f,\Phi):=\displaystyle\int_{\mathbb{V}(\mathbb{A})}f(\xi_{0}\cdot v\cdot\iota_{SL_{2}}\cdot\widetilde{h})\overline{\omega_{S}(v(\iota_{SL_{2}}\cdot\widetilde{h}))\Phi(0)}dv$ is a section of $\widetilde{I}_{1}(\omega\chi_{S},s)$ where $\chi_{S}(\ast)=\langle-S^{n},\ast\rangle_{\mathbb{A}}=\langle-S,\ast\rangle_{\mathbb{A}}$ is defined by the Hilbert symbol $\langle\ast,\ast\rangle_{\mathbb{A}}$, and
  3. (3)
    $E_{\psi_{S},\Phi}(\ast;f)$ is an Eisenstein series on ${\rm SL}_{2}(\mathbb{A})$ or $\widetilde{{\rm SL}_{2}(\mathbb{A})}$ associated to $R(\ast;f,\Phi)$ according to ${\rm dim}\hskip 0.56905ptX=n$ is even or odd.
7. Fourier-Jacobi expansions of $E_{l}(g,\Phi_{\mathbb{f}})$ along $J$: an explicit form
Recall the Eisenstein series $E_{l}(g,\Phi_{f}):=E_{l}(g,\Phi_{\mathbb{f}},l+1)$ defined by the Pollack’s section $f_{l}(\ast,\Phi_{\mathbb{f}},l+1)\in I(\mathbf{1},l-\frac{n}{2})={\rm Ind}^{G(\mathbb{A})}_{P(\mathbb{A})}|\nu|^{l-\frac{n+2}{2}}$ (see Section 4). In this section, we explicitly compute the Fourier-Jacobi coefficients of $E_{l}(g,\Phi_{\mathbb{f}})$.
7.1. Preliminaries at the archimedean place
For each $a,S\in\mathbb{Q}$, put $\eta:=(a,\overbrace{0,\ldots,0}^{n},S)\in V^{\prime}(\mathbb{Q})$ and we view it as
$\eta=(0,a,\overbrace{0,\ldots,0}^{n},S,0)\in V(\mathbb{Q}).$
We assume that
(7.1) $(\eta,\eta)_{V^{\prime}}:=\frac{1}{2}\eta J_{2,n}{}^{t}\eta=aS>0.$
Hence $a$ and $S$ have the same sign. Notice that the action of $m\in{\rm SO}(V)$ on $\textbf{x}\in V$ is defined by
$m\textbf{x}:=\textbf{x}\cdot{}^{t}m$
as a matrix multiplication. Further, for any $\textbf{x},\ \textbf{y}\in V$ and $m\in{\rm SO}(V)$,
$(\textbf{x},m\textbf{y})_{V}=\frac{1}{2}\textbf{x}J({}^{t}(\textbf{y}\cdot{}^{t}m))=\frac{1}{2}\textbf{x}(Jm){}^{t}\textbf{y}=\frac{1}{2}\textbf{x}({}^{t}m^{-1}J){}^{t}\textbf{y}=(m^{-1}\textbf{x},\textbf{y})_{V}.$
In the computation later, we need to consider
(7.2) $v(x,0,0)\ell({\rm diag}(t,t^{-1}))={\rm diag}(t,m,t^{-1})\in M(\mathbb{R}),m:=v(x,0,0)\ell({\rm diag}(1,t^{-1}))$
for $t\in GL_{1}$ and $x=(x_{1},x^{\prime},x_{n})\in X$ where $x_{1},x_{n}\in\mathbb{G}_{a}$ and $x^{\prime}\in\mathbb{G}^{n-2}_{a}$. For such $(t,m)\in M(\mathbb{R})$, we compute
(7.3) $u_{2\pi\eta}(t,m)$ $=$ $it(2\pi\eta,m(iv_{1}-v_{2}))$
$=$ $-2\pi i\{(at^{2}+S)-i\cdot St(x_{1}+x_{n})-St^{2}x_{1}x_{n}+St^{2}(x^{\prime},x^{\prime})\}.$
where $(x^{\prime},x^{\prime})=\frac{1}{2}x^{\prime}A{}^{t}x^{\prime}$ and $i=\sqrt{-1}$.
For above $(t,m)\in M(\mathbb{R})$ with $m:=v(x,0,0)\ell({\rm diag}(1,t^{-1}))$ in (7.2), we will also compute
(7.4) $I_{\eta}(t,\alpha,\beta;\Phi):=\int_{X(\mathbb{R})}\mathcal{W}_{2\pi\eta}(t,m)\overline{F_{\alpha,\beta}(x;\Phi)}dx,\ t,\alpha,\beta\in\mathbb{R},\ t>0,\beta>0,\ \Phi\in\mathcal{S}(X(\mathbb{R}))$
where
(7.5) $F_{\alpha,\beta}(x;\Phi)$ $=$ $\sqrt{\beta}^{\frac{n}{2}}\Phi(\sqrt{\beta}x)e^{2\pi\sqrt{-1}S\frac{\alpha}{2}\sigma(x,x)},\ \Phi\in S(X(\mathbb{R}))$
and
$\mathcal{W}_{2\pi\eta}(t,m)=t^{l+1}\sum_{v=-l}^{l}\Bigg(\frac{|u_{2\pi\eta}(t,m)|}{u_{2\pi\eta}(t,m)}\Bigg)^{v}K_{v}(|u_{2\pi\eta}(t,m)|)X^{l-v}Y^{l+v}$
where $u_{2\pi\eta}(t,m)$ is given explicitly in (7.3). Here the exponent $\frac{n}{2}=\frac{1}{2}{\rm dim}X$ in $F_{\alpha,\beta}(x;\Phi)$ comes from the action of the Weil representation. Note that by using $K_{v}=K_{-v}$, we can also write it as
$\mathcal{W}_{2\pi\eta}(t,m)=t^{l+1}\sum_{v=-l}^{l}\Bigg(\frac{u_{2\pi\eta}(t,m)}{|u_{2\pi\eta}(t,m)|}\Bigg)^{v}K_{v}(|u_{2\pi\eta}(t,m)|)X^{l+v}Y^{l-v}.$
We write (7.4) as
(7.6) $I_{\eta}(t,\alpha,\beta;\Phi)=\sum_{v=-l}^{l}I_{\eta}(t,\alpha,\beta;\Phi)_{v}X^{l+v}Y^{l-v}$
where
(7.7) $I_{\eta}(t,\alpha,\beta;\Phi)_{v}:=\int_{X(\mathbb{R})}t^{l+1}\Bigg(\frac{u_{2\pi\eta}(t,m)}{|u_{2\pi\eta}(t,m)|}\Bigg)^{v}K_{v}(|u_{2\pi\eta}(t,m)|)\overline{F_{\alpha,\beta}(x;\Phi)}dx$
for above $(t,m)\in M(\mathbb{R})$.
Changing $(x_{1},x_{n})$ with $(t_{1}+t_{n},t_{1}-t_{n})$ with the Jacobian $2$, the equation (7.3) becomes
(7.8) $u_{2\pi\eta}(t,m)$ $=$ $i\cdot{\rm sgn}(S)\{(\lambda\cdot t_{1}+t^{-1}\lambda\cdot i)^{2}-\mu\}$
where
$\lambda:=t\sqrt{2\pi|S|},\ \mu:=2\pi t^{2}\Big(|a|+|S|t^{2}_{n}+|S|(x^{\prime},x^{\prime})\Big).$
Henceforth we assume $t=\sqrt{\beta}$ for $\beta>0$ so that
$\lambda=\sqrt{\beta}\sqrt{2\pi|S|},\ \mu=2\pi\beta\Big(|a|+|S|t^{2}_{n}+|S|(x^{\prime},x^{\prime})\Big)$
Note that $\lambda,\mu>0$ since $aS,\beta>0$ and $(x^{\prime},x^{\prime})\geq 0$. For each $S\in\mathbb{Q}^{\times}$ and $r\in\mathbb{Z}_{\geq 0}$, put
(7.9) $\Phi^{1}_{S,r}(t_{1})=(2\pi|S|)^{\frac{r}{2}}t^{r}_{1}e^{-2\pi|S|t^{2}_{1}}$
so that $\Phi^{1}_{S,r}(\sqrt{\beta}t_{1})=(\lambda t_{1})^{r}e^{-\lambda^{2}t^{2}_{1}}$ and
(7.10) $\Phi_{S,r}:=\Phi^{1}_{S,r}\otimes\Phi^{\prime}\in S(X(\mathbb{R}))$
where $\Phi^{\prime}\in S(\mathbb{R}^{n-1})$ with respect to the coordinate $(x^{\prime},t_{n})\in\mathbb{R}^{n-1}\subset V^{\prime\prime}(\mathbb{R})=X(\mathbb{R})$. We remark that the restriction of the quadratic space $V^{\prime\prime}$ to $\mathbb{R}^{n-1}$ is negative definite. Applying Theorem 12.1 and Remark 12.2 to the case when $v=0$, we have that
$I_{\eta}(\sqrt{\beta},0,\beta,\Phi_{S,r})_{0}=2\beta^{\frac{l+1}{2}+\frac{n}{4}}\times$
$\int_{(x^{\prime},t_{n})\in\mathbb{R}^{n-1}}\Bigg(\int_{t_{1}\in\mathbb{R}}(\lambda t_{1})^{r}e^{-\lambda^{2}t^{2}_{1}}K_{0}(|(\lambda\cdot t_{1}+\sqrt{\beta}^{-1}\lambda\cdot i)^{2}-\mu|)dt_{1}\Bigg)\overline{\Phi^{\prime}(\sqrt{\beta}x^{\prime},\sqrt{\beta}t_{n})}dx^{\prime}dt_{n}$
$=$ $2\beta^{\frac{l+1}{2}+\frac{n}{4}}\lambda^{-1}C(S,r)\Bigg(\int_{(x^{\prime},t_{n})\in\mathbb{R}^{n-1}}e^{-\mu}\overline{\Phi^{\prime}(\sqrt{\beta}x^{\prime},\sqrt{\beta}t_{n})}dx^{\prime}dt_{n}\Bigg)$
$=$ $2\beta^{\frac{l+1}{2}+\frac{n}{4}}\lambda^{-1}C(S,r)C(S,\beta,\Phi^{\prime})e^{-2\pi\beta|a|}$
where
(7.11) $C(S,r):=\left\{\begin{array}[]{ll}0&\text{if $r$ is odd,}\\ C(r,\sqrt{\beta}^{-1}\lambda)=\displaystyle\int_{t\in\mathbb{R}}t^{r}e^{-t^{2}}K_{0}(t^{2}+2\pi|S|)dt&\text{if $r$ is even,}\end{array}\right$.
is the constant given in Theorem 12.1 for $r$ even (we use the notation $C(r,\sqrt{\beta}^{-1}\lambda)$ therein) and
$C(S,\beta,\Phi^{\prime}):=\int_{(x^{\prime},t_{n})\in\mathbb{R}^{n-1}}e^{-2\pi\beta|S|\{t_{n}^{2}+(x^{\prime},x^{\prime})\}}\overline{\Phi^{\prime}(\sqrt{\beta}x^{\prime},\sqrt{\beta}t_{n})}dx^{\prime}d_{t_{n}}.$
By the change of variables, we have
$C(S,\beta,\Phi^{\prime})=C_{1}(S,\Phi^{\prime})\beta^{-\frac{n-1}{2}},\ C_{1}(S,\Phi^{\prime}):=C(S,1,\Phi^{\prime}).$
Thus, we have
(7.12) $I_{\eta}(\sqrt{\beta},0,\beta,\Phi_{S,r})_{0}=C_{2}(S,r)\sqrt{\beta}^{l-\frac{n}{2}+1}e^{-2\pi|a|\beta}$
where $C_{2}(S,\Phi_{S,r})=\displaystyle\frac{2C(S,r)C_{1}(S,\Phi^{\prime})}{\sqrt{2\pi|S|}}$.
Remark 7.1.
Let $S\in\mathbb{Q}^{\times}$. If $\Phi_{S,0}$ is Gaussian, the constant $C_{2}(S,\Phi_{S,0})$ is non-zero.
Since $\{\Phi^{1}_{S,r}\}_{r\geq 0}$ is a topological basis of $\mathcal{S}(\mathbb{R})$, we have the following result:
Theorem 7.2.
Let $S\in\mathbb{Q}^{\times}$. For any $\Phi\in\mathcal{S}(X(\mathbb{R}))$, there exists a constant $C_{3}(S,\Phi)$ depending only on $S$ and $\Phi$ such that
$I_{\eta}(\sqrt{\beta},0,\beta,\Phi_{S})_{0}=C_{3}(S,\Phi)\sqrt{\beta}^{l-\frac{n}{2}+1}e^{-2\pi|a|\beta}.$
Similarly, for any $v$-th component, we have
Theorem 7.3.
Let $S\in\mathbb{Q}^{\times}$. For even $r\geq 0$ and $v$ with $l\leq v\leq l$,
$I_{\eta}(\sqrt{\beta},0,\beta,\Phi_{S,r})_{v}=C_{2}(S,\Phi_{S,r})(-i)^{v}{\rm sgn}(S)^{v}\sqrt{\beta}^{l-\frac{n}{2}+1}e^{-2\pi|a|\beta}.$
7.2. Computation of Fourier-Jacobi coefficients
Now, we are ready to compute the Fourier-Jacobi expansion of $E_{l}$ when $s=l+1$. Recall the settings in Section 5 and (4.1). Let $w_{1}=\left(\begin{array}[]{ccc}0&0&1\\ 0&I_{n}&0\\ 1&0&0\end{array}\right)\in{\rm SO}(V^{\prime})(\mathbb{Q})$ so that
$w_{1}(x_{1},x,x_{n})w^{-1}_{1}=(x_{n},x,x_{1}),\ (x_{1},x,x_{n})\in V^{\prime}.$
By using this and automorphy in $w_{1}$, changing the coordinates of ${\rm SO}(V^{\prime})$ by $w_{1}$ (hence ${\rm SO}(V^{\prime})\ni m\mapsto w^{-1}_{1}m\in{\rm SO}(V^{\prime})$), and replacing $\Phi$ with $w_{1}\cdot\Phi$, it is easy to see that for $n(x){\rm diag}(t,m,t^{-1})\in N(\mathbb{A})M(\mathbb{A})=P(\mathbb{A})$,
$E_{l}(n(x)(t,m),\Phi_{\mathbb{f}},l+1)=E_{0}(n(x)(t,w^{-1}_{1}m,t^{-1}))$
(7.13) $\phantom{xxxxxssssssssssx}+\sum_{\eta\in V(\mathbb{Q})\atop q(\eta)\geq 0}a_{E_{l}}(\eta)(n(x_{\mathbb{f}})(t_{\mathbb{f}},m_{\mathbb{f}},t^{-1}_{\mathbb{f}})\Phi_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(t_{\infty},m_{\infty})\psi_{w^{-1}_{1}\eta w_{1}}(n(x))$
where $\psi_{w^{-1}_{1}\eta w_{1}}(n(x))=\psi((w^{-1}_{1}\eta w_{1},x)_{V^{\prime}})$ and $\psi:\mathbb{A}\longrightarrow\mathbb{C}^{\times}$ is the standard additive character.
Let $S\in\mathbb{Q}^{\times}$ and $\Phi=\Phi_{\mathbb{f}}\otimes\Phi_{\infty}\in S(X(\mathbb{A}))$. We define the Fourier-Jacobi coefficient of $E_{l}(\ast,\Phi_{\mathbb{f}},l+1)$ with respect to $\Phi$ and of index $S$ by
[…]
where $\widetilde{h}=(h,\varepsilon)\in\widetilde{H^{\rm ss}(\mathbb{A})}$. Notice that $n(x)=v(0,0,z)$ for $x=(\overbrace{0,\ldots,0}^{n+1},z)\in V^{\prime}$ and for such $x$ and $\eta=(\eta_{1},\ldots,\eta_{n+2})\in V^{\prime}(\mathbb{Q})$, we have
$\psi_{w^{-1}_{1}\eta w_{1}}(n(x))=\psi(\tfrac{1}{2}\eta_{n+2}z).$
Thus, it is easy to see that
[…]
where
$E_{l}(g,\Phi_{\mathbb{f}},l+1)_{S}=\sum_{\eta=(\ast,\ldots,\ast,2S)\in V^{\prime}(\mathbb{Q})\atop q(\eta)\geq 0}a_{E_{l}}(\eta)(g_{\mathbb{f}}\cdot\Phi_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A}).$
It is easy to see that $v(x,y,z)=v(0,y,z+(x,y)_{V^{\prime\prime}})v(x,0,0)$ where $(x,y)_{V^{\prime\prime}}=\frac{1}{2}xJ_{1,n-1}{}^{t}y=\frac{1}{2}\sigma(x,y)$. Then, for $\eta=(\eta_{1},\eta^{\prime},2S)\in V^{\prime}(\mathbb{Q})$ with $\eta^{\prime}\in V^{\prime\prime}(\mathbb{Q}),\ \eta_{1}\in\mathbb{Q}$,
$\psi_{w^{-1}_{1}\eta w_{1}}(v(0,y,z+(x,y)_{V^{\prime\prime}}))=\psi(\tfrac{1}{2}\sigma(\eta^{\prime},y))\psi(S(z+\tfrac{1}{2}\sigma(x,y))).$
Then, we have
[…]
(7.15) $(a,2S\xi,2S)=v(-\xi,0,0)\cdot\eta_{a,S,\xi}\cdot v(\xi,0,0),\ \eta_{a,S,\xi}=(a+S\sigma(\xi,\xi),\overbrace{0,\ldots,0}^{n},S).$
[…]
where
$E_{\xi,S,\Phi}(\widetilde{h})=\sum_{a\in\mathbb{Q}\atop aS+\frac{S^{2}}{2}\sigma(\xi,\xi)\geq 0}\int_{X(\mathbb{A}_{f})}a_{E_{l}}(\eta_{a,S,\xi})(v(x_{\mathbb{f}},0,0)\widetilde{h}_{\mathbb{f}}\cdot\Phi_{\mathbb{f}})\overline{(\omega_{\psi_{S,\mathbb{f}}}(v(x_{\mathbb{f}},0,0)\widetilde{h}_{\mathbb{f}})\Phi_{\mathbb{f}})(0)}dx_{\mathbb{f}}$
$\times\int_{X(\mathbb{R})}\mathcal{W}_{2\pi\eta_{a,S,\xi}}(v(x_{\infty},0,0)\widetilde{h}_{\infty})\overline{(\omega_{\psi_{S,\infty}}(v(x_{\infty},0,0)\widetilde{h}_{\infty})\Phi_{\infty})(0)}dx_{\infty}.$
Note that (7.16) is a finite sum (see [15, Section 5.2]).
Thus we have
$E_{\xi,S,\Phi}(\widetilde{h})=\sum_{a\in\mathbb{Q}\atop aS+\frac{S^{2}}{2}\sigma(\xi,\xi)\geq 0}a_{E_{l}}(\eta_{a,S,\xi})(\widetilde{h}_{\mathbb{f}}\Phi_{\mathbb{f}})I_{\eta}(a,S,\xi;\Phi_{\infty})(\widetilde{h}_{\infty}),\ \widetilde{h}=\widetilde{h}_{\mathbb{f}}\widetilde{h}_{\infty}\in H^{\rm ss}(\mathbb{A})$
where
(7.17) $I_{\eta}(a,S,\xi;\Phi_{\infty})(\widetilde{h}_{\infty}):=\int_{X(\mathbb{R})}\mathcal{W}_{2\pi\eta_{a,S,\xi}}(v(x_{\infty},0,0)\widetilde{h}_{\infty})\overline{(\omega_{\psi_{S,\infty}}(v(x_{\infty},0,0)\widetilde{h}_{\infty})\Phi_{\infty})(0)}dx_{\infty}.$
[…]
$v(x_{\infty},0,0)h_{\infty}=\Bigg(v(x_{\infty},0,0)\begin{pmatrix}1&\alpha\\ 0&1\end{pmatrix}v(-x_{\infty},0,0)\Bigg)v(x_{\infty},0,0){\rm diag}(\sqrt{\beta},\sqrt{\beta}^{-1})$
$n(-\alpha,\alpha x_{\infty},\frac{\alpha}{2}\sigma(x_{\infty},x_{\infty}))v(x_{\infty},0,0){\rm diag}(\sqrt{\beta},\sqrt{\beta}^{-1})\in N(\mathbb{R})M(\mathbb{R})$
and
$v(x_{\infty},0,0){\rm diag}(\sqrt{\beta},\sqrt{\beta}^{-1})={\rm diag}(\sqrt{\beta},m,\sqrt{\beta}^{-1}),\ m=v(x_{\infty},0,0)\ell(1,\sqrt{\beta}^{-1}).$
Using this, we have
$\mathcal{W}_{2\pi\eta_{a,S,\xi}}(v(x_{\infty},0,0)h_{\infty})=\psi_{w_{1}(2\pi\eta_{a,S,\xi})w^{-1}_{1}}(n(-\alpha,\alpha x_{\infty},\tfrac{\alpha}{2}\sigma(x_{\infty},x_{\infty})))$
$\phantom{xxxxxxxxxxxxxxx}\times\mathcal{W}_{2\pi\eta_{a,S,\xi}}(v(x_{\infty},0,0){\rm diag}(\sqrt{\beta},\sqrt{\beta}^{-1}))$
$\psi_{\infty}(-\alpha(a+S\sigma(\xi,\xi)))\psi_{\infty}(\tfrac{S\alpha}{2}\sigma(x_{\infty},x_{\infty}))\mathcal{W}_{2\pi\eta_{a,S,\xi}}(\beta,m)$
while by definition of the Weil representation and recalling (7.5),
$\omega_{\psi_{S,\infty}}(v(x_{\infty},0,0)\widetilde{h}_{\infty})\Phi_{\infty})(0)=\varepsilon^{n}_{\infty}\frac{\gamma_{\infty}(1)}{\gamma_{\infty}(S^{n})}F_{\alpha,\beta}(x_{\infty};\Phi_{\infty})$
$\phantom{xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx}=\varepsilon^{n}_{\infty}\frac{\gamma_{\infty}(1)}{\gamma_{\infty}(S^{n})}\beta^{\frac{n}{4}}\Phi_{\infty}(\sqrt{\beta}x_{\infty})\psi_{\infty}(\tfrac{S\alpha}{2}\sigma(x_{\infty},x_{\infty})).$
Therefore, we have
$I_{\eta}(a,S,\xi;\Phi_{\infty})(h_{\infty})=\varepsilon^{n}_{\infty}\frac{\gamma_{\infty}(1)}{\gamma_{\infty}(S^{n})}e^{-2\pi\alpha\sqrt{-1}(a+S\sigma(\xi,\xi))}\int_{X(\mathbb{R})}\mathcal{W}_{2\pi\eta_{a,S,\xi}}(\beta,m)\overline{F_{0,\beta}(x;\Phi_{\infty})}dx_{\infty}$
$\varepsilon^{n}_{\infty}\frac{\gamma_{\infty}(1)}{\gamma_{\infty}(S^{n})}e^{-2\pi\alpha\sqrt{-1}(a+S\sigma(\xi,\xi))}I_{\eta_{a,S,\xi}}(\sqrt{\beta},0,\beta,\Phi_{\infty})$
in terms of the notation (7.4). Summing up, we have
Proposition 7.4.
Keep the notation being as above. It holds that
(7.18) $E_{\xi,S,\Phi}(\widetilde{h})=\varepsilon^{n}_{\infty}\frac{\gamma_{\infty}(1)}{\gamma_{\infty}(S^{n})}\sum_{a\in\mathbb{Q}\atop aS+S^{2}\sigma(\xi,\xi)\geq 0}a_{E_{l}}(\eta_{a,S,\xi})(h_{\mathbb{f}}\Phi_{\mathbb{f}})I_{\eta_{a,S,\xi}}(\sqrt{\beta},0,\beta,\Phi_{\infty})e^{-2\pi\alpha\sqrt{-1}(a+S\sigma(\xi,\xi))}$
for $\widetilde{h}=(h,\varepsilon)\in\widetilde{{\rm SL}_{2}(\mathbb{A})}$ with $h_{\infty}\sqrt{-1}=\alpha+\beta\sqrt{-1}$.
According to (7.6), we write
(7.19) $E_{\xi,S,\Phi}(\widetilde{h})=\sum_{v=-l}^{l}E_{\xi,S,\Phi}(\widetilde{h})_{v}X^{l-v}Y^{l+v}$
[…]
Corollary 7.5.
Keep the notation being as above. Let $\Phi=\Phi_{\mathbb{f}}\otimes\Phi_{\infty}\in\mathcal{S}(X(\mathbb{A}))$. If $v\neq 0$, suppose that $\Phi_{\infty}=\Phi_{S,r}$ with even $r\geq 0$. Then, if $S>0$,
$E_{\xi,S,\Phi}(\widetilde{h})_{v}=C_{4}(S,\Phi_{\infty})\varepsilon^{n}_{\infty}\sqrt{\beta}^{l-\frac{n}{2}+1}\sum_{a\in\mathbb{Q}_{>0}\atop a+S\sigma(\xi,\xi)\geq 0}a_{E_{l}}(\eta_{a,S,\xi})(h_{\mathbb{f}}\Phi_{\mathbb{f}})(\bar{q}_{\tau})^{a+S\sigma(\xi,\xi)}$
and if $S<0$,
$E_{\xi,S,\Phi}(\widetilde{h})_{0}=C_{4}(S,\Phi_{\infty})\varepsilon^{n}_{\infty}\sqrt{\beta}^{l-\frac{n}{2}+1}\sum_{a\in\mathbb{Q}_{<0}\atop a+S\sigma(\xi,\xi)<0}a_{E_{l}}(\eta_{a,S,\xi})(h_{\mathbb{f}}\Phi_{\mathbb{f}})(q_{\tau})^{-(a+S\sigma(\xi,\xi))}$
where $C_{4}(S,\Phi_{\infty})=\frac{\gamma_{\infty}(1)}{\gamma_{\infty}(S^{n})}\times\left\{\begin{array}[]{ll}C_{3}(S,\Phi_{\infty})&\text{if $v=0$}\\ C_{2}(S,\Phi_{S,r})&\text{if $v\neq 0$}\end{array}\right$. and $q_{\tau}=e^{2\pi\sqrt{-1}\tau}$ with $\tau=\alpha+\beta\sqrt{-1}$.
Remark 7.6.
For each $\Phi=\Phi_{\mathbb{f}}\otimes\Phi_{\infty}\in\mathcal{S}(X(\mathbb{A}))$ and $\xi\in X(\mathbb{R})$, there exist $\Phi^{(1)}_{\xi,\mathbb{f}},\ldots,\Phi^{(r)}_{\xi,\mathbb{f}}\in\mathcal{S}(X(\mathbb{A}_{\mathbb{f}}))$ and constants $c_{1},\ldots,c_{r}\in\mathbb{C}$ such that
$E_{\xi,S,\Phi}=\sum_{i=1}^{r}c_{i}E_{S,\Phi^{(i)}_{\xi,\mathbb{f}}\otimes\Phi_{\infty}}$
(this follows from the argument for [15, Section 5.2] and the proof of [16, Lemma 5.4]).
The same is true for each $v$-th component $(E_{\xi,S,\Phi})_{v}$.
We denote by $\kappa_{\theta}\in\widetilde{{\rm SO}(2)}$ corresponding to $e^{i\theta}\in U(1)$ under the natural identification $\widetilde{{\rm SO}(2)}\simeq{\rm SO}(2)\simeq U(1)$. Let $\Phi_{\infty}\in\mathcal{S}(X(\mathbb{R}))$ be an $\widetilde{{\rm SO}(2)}-$eigen function of $S(X(\mathbb{R}))$ for $\omega_{S,\infty}$. It follows from (5.1) that there exists an integer $k\in\mathbb{Z}_{\geq 0}$ such that
(7.20) $\omega_{S,\infty}(\kappa_{\theta})\Phi_{\infty}=e^{i\theta\cdot{\rm sgn}(S)(\frac{2-n}{2}+k)}\Phi_{\infty},\ \kappa_{\theta}\in\widetilde{{\rm SO}(2)}$.
Thus, $\Phi_{\infty}$ is of $\widetilde{{\rm SO}(2)}-$weight ${\rm sgn}(S)(\frac{2-n}{2}+k)$. The bottom case of $k=0$ is attained by the Gaussian on $X(\mathbb{R})$.
The following result is an immediate consequence of Theorem 6.1 with Remark 7.6 and it is consistent with Corollary 7.5:
Corollary 7.7.
Keep the notation being as above. Let $\Phi=\Phi_{\mathbb{f}}\otimes\Phi_{\infty}\in\mathcal{S}(X(\mathbb{A}))$ and assume $\Phi_{\infty}$ is of $\widetilde{{\rm SO}(2)}-$weight (-type) ${\rm sgn}(S)(\frac{2-n}{2}+k)$. Then, the $v$-th component $(E_{\xi,S,\Phi})_{v}$ for each $l\leq v\leq l$ is the Eisenstein series defined by a section of $\left\{\begin{array}[]{cc}I_{1}(\mathbf{1},l-\frac{n}{2})&\text{if $n$ is even}\\ \widetilde{I}_{1}(\chi_{S},l-\frac{n}{2})&\text{if $n$ is odd}\end{array}\right$. whose $\left\{\begin{array}[]{cc}\text{${\rm SO}(2)$-type}&\text{if $n$ is even}\\ \text{$\widetilde{{\rm SO}(2)}$-type}&\text{if $n$ is odd}\end{array}\right$. is $v-{\rm sgn}(S)\Big(\frac{2-n}{2}+k\Big).$ In particular, $E_{\xi,S,\Phi}(\widetilde{h})_{v}$ corresponds to
[…]
Remark 7.8.
For any modular form $f:\mathbb{H}\longrightarrow\mathbb{C}$ of weight $k(\in\mathbb{Z}_{\geq 0})$ with respect to a suitable congruence subgroup, one can associate an automorphic form $\varphi_{f}$ on ${\rm SL}_{2}(\mathbb{A})$ so that $\varphi_{f}((1_{\mathbb{f}},h_{\infty}))=j(h_{\infty},\sqrt{-1})^{k}f(h_{\infty}\sqrt{-1})$. Conversely, any automorphic form $\varphi$ on ${\rm SL}_{2}(\mathbb{A})$ for which $j(h_{\infty},\sqrt{-1})^{-k}\varphi((h_{\mathbb{f}},h_{\infty}))$ is holomorphic for every fixed $h_{\mathbb{f}}\in{\rm SL}_{2}(\mathbb{A}_{\mathbb{f}})$ arises from a modular form of weight $k$ in the above manner. The last statement of Corollary 7.7 follows this sense.
8. A uniform proof for the main theorems
In this section, we prove Theorem 1.1 and 1.2. We follow the method in [10], [11], [15] rather than representation theoretic arguments [14], [16].
8.1. A formal series
Recall the Laurent polynomial $\widetilde{Q}_{\eta,p}(X_{p})$ depending on $\eta\in V^{\prime}(\mathbb{Q}_{p})$ and $\Phi_{p}={\rm char}_{V^{\prime}(\mathbb{Z}_{p})}$. We write it as $\widetilde{Q}_{\eta,\Phi_{p}}(X_{p}):=\widetilde{Q}_{\eta,p}(X_{p})$ to indicate independence. By Lemma 4.1 and the computation in Section 4, for each $g_{p}={\rm diag}(t_{p},m_{p},t_{p}^{-1})n_{p}\in P(\mathbb{Q}_{p})=M(\mathbb{Q}_{p})N(\mathbb{Q}_{p})$, we see that
(8.1) $\widetilde{Q}_{\eta,g^{-1}_{p}\Phi_{p}}(X_{p})=|t_{p}|^{\frac{n-2}{2}}_{p}\widetilde{Q}_{g_{p}^{-1}\eta,\Phi_{p}}(X_{p})$
(notice that $q(g_{p}^{-1}\eta)=t^{-2}_{p}q(\eta)$ and $\chi_{\eta}=\chi_{g_{p}^{-1}\eta}$).
8.1.1. The even case
Assume $n$ is even. Put $k=l-\frac{n-2}{2}$. For each $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$ and $g_{\mathbb{f}}=(g_{p})_{p}\in G(\mathbb{A}_{\mathbb{f}})$, we define
$A(\eta;\{X_{p}\}_{p})(g_{\mathbb{f}}):=q(\eta)^{\frac{k-1}{2}}\prod_{p}\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})\in\otimes^{\prime}_{p<\infty}\mathbb{C}[X_{p},X^{-1}_{p}].$
This is well-defined since $\widetilde{Q}_{\eta,p}=1$ if $p\nmid q(\eta)$. Further, it is zero unless $\eta\in V^{\prime}(\mathbb{Z})$ since $J(s,\eta,f_{l})=0$ otherwise. We also define
(8.2) $F(g;\{X_{p}\}_{p}):=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A(\eta;\{X_{p}\}_{p})(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\quad g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A}).$
8.1.2. The odd case
Assume $n$ is odd. Put k=2ln+1k=2l-n+1. Recall the notation right before Theorem 1.2. For each $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$ and $g_{\mathbb{f}}=(g_{p})_{p}\in G(\mathbb{A}_{\mathbb{f}})$, we define
$A(\eta;\{X_{p}\}_{p})(g_{\mathbb{f}}):=c(\mathfrak{d}_{\eta})\mathfrak{f}_{\eta}^{\frac{k-1}{2}}\prod_{p}\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})\in\otimes^{\prime}_{p<\infty}\mathbb{C}[X_{p},X^{-1}_{p}].$
We also define
(8.3) $F(g;\{X_{p}\}_{p}):=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A(\eta;\{X_{p}\}_{p})(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\ g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A}).$
8.2. A compatible family of Eisenstein series
Let $l^{\prime}$ be a sufficiently large positive even integer greater than $n+1$. Let $k^{\prime}=l^{\prime}-\frac{n}{2}+1$. By the results in Sections 4 and 7, we see that
$F(g;\{p^{l^{\prime}-\frac{n}{2}}\}_{p})=\frac{1}{C}E_{l^{\prime}}(g,\Phi_{\mathbb{f}})$
where $\Phi_{\mathbb{f}}={\rm char}_{X(\widehat{\mathbb{Z}})}$ as considered in Section 4 and the constant $C=C_{l^{\prime},n}$ or $C^{\prime}_{l^{\prime},n}$ is given in Theorem 4.8 and Theorem 4.9 respectively.
We now assume $S<0$. Let $\Phi_{\mathbb{f}}={\rm char}_{X(\widehat{\mathbb{Z}})}$ as above. For $v=0$ or $l^{\prime}$, we define $\Phi^{(l^{\prime})}_{\infty}$ to be any element in $\mathcal{S}(X(\mathbb{R}))$ of weight $k^{\prime}$ if $v=0$ and the Gaussian $\Phi_{S,0}$ if $v=l^{\prime}$. Let $\Phi^{(l^{\prime})}=\Phi_{\mathbb{f}}\Phi^{(l^{\prime})}_{\infty}$. By Corollary 7.7, for $v=0,l^{\prime}$, the $v$-th component $(E_{\xi,S,\Phi^{(l^{\prime})}})_{v}$ corresponds to a holomorphic modular form of weight $k^{\prime}$ in the sense of Remark 7.8. Further, by Corollary 7.5,
[…]
$=\varepsilon^{n}_{\infty}C_{4}(S,\Phi^{(l^{\prime})}_{\infty})\sum_{a\in\mathbb{Q}_{<0}\atop a+S\sigma(\xi,\xi)\leq 0}(a+S\sigma(\xi,\xi))^{\frac{k^{\prime}-1}{2}}A(\eta_{a,S,\xi};\{p^{l^{\prime}-\frac{n}{2}}\})(h_{\mathbb{f}})(q_{\tau})^{-(a+S\sigma(\xi,\xi))},$
where $j_{\frac{1}{2}}$ stands for the automorphic factor of weight $\frac{1}{2}$. Thus,
$\{S^{-\frac{k^{\prime}-1}{2}}j_{\frac{1}{2}}(h_{\infty},\sqrt{-1})^{2k^{\prime}}(E_{\xi,S,\Phi^{(l^{\prime})}})_{v}(\widetilde{h})\}_{l^{\prime}\gg n+1,\xi\in X(\mathbb{Q})}$
makes up a compatible family of Eisenstein series in the sense of [11, Section 9] for $n$ even and of [10, Section 10] for $n$ odd.
The case when $S>0$ is similarly handled and we also have a compatible family of Eisenstein series for $v=0,-l^{\prime}$. In fact, we have a family of anti-holomorphic Eisenstein series but the above mentioned results also holds in this setting.
8.3. Automorphy of the formal series
Let $f=\displaystyle\sum_{m=1}^{\infty}a_{f}(m)m^{\frac{k-1}{2}}q^{m}$ be a normalized Hecke eigen cusp form of weight $k=\left\{\begin{array}[]{ll}l-\frac{n}{2}+1&\text{for $n$ even,}\\ 2l-n+1&\text{for $n$ odd}\end{array}\right$. with respect to ${\rm SL}_{2}(\mathbb{Z})$. For each prime $p$, let $a_{f}(p)=\alpha_{p}+\alpha^{-1}_{p}$. We will prove that the formal series
$F_{f}(g):=F(g;\{\alpha_{p}\}_{p})=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A(\eta;\{\alpha_{p}\}_{p})(g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(g_{\infty}),\quad g=g_{\mathbb{f}}g_{\infty}\in G(\mathbb{A}),$
is a cuspidal automorphic form of level one in the sense of Section 3.
First we check the non-vanishing of $F_{f}$:
Lemma 8.1.
There exists $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$ such that $A(\eta;\{\alpha_{p}\}_{p})\neq 0$.
Proof.
When $n$ is even, it is easy to find $\eta$ with $q(\eta)=1$. Thus, we have $A(\eta;\{\alpha_{p}\}_{p})=1$ by definition.
Now assume $n$ is odd. Let $D$ be a square-free fundamental discriminant. Then for $\eta=(D,D,\overbrace{0,\ldots,0}^{n},\pm 1,\pm 1)$, $q(\eta)=\pm D$, where we choose a sign so that $\pm D>0$. Then by Theorem 4.3, Theorem 4.9, and (8.3), $A(\eta;\{\alpha_{p}\}_{p})=c(|D|)$. By Kohnen-Zagier [18, Theorem 1], $c(|D|)$ is proportional to the central value of $f\otimes\chi_{D}$. The claim follows from [22, p.654]. ∎
Henceforth, we assume that $n$ is even. The odd case is similarly proved.
First, we prove the convergence:
Lemma 8.2.
The formal series $F_{f}(g)$ is absolutely convergent and of moderate growth in $g_{\infty}$.
Proof.
We first show
$\left|\prod_{p|q(\eta)}\widetilde{Q}_{\eta,p}(\alpha_{p})\right|\ll q(\eta)^{\frac{n}{2}+\epsilon}.$
Case 1: $n$ even: Note the formula for $\widetilde{Q}_{\eta,p}$ from (11.5). Then since $|\alpha_{p}|\leq 1$,
$|\widetilde{Q}_{\eta,p}(\alpha_{p})|\leq 2\sum_{a=0}^{k-1}A_{a}+(k^{\prime}+1)A_{k}.$
Using the crude estimate $k^{\prime}+1\leq p^{k^{\prime}}$, we have $|\widetilde{Q}_{\eta,p}(\alpha_{p})|\leq 4p^{\frac{n}{2}(2k+k^{\prime})}$. Hence
$\left|\prod_{p|q(\eta)}\widetilde{Q}_{\eta,p}(\alpha_{p})\right|\leq 4^{\omega(q(\eta)}q(\eta)^{\frac{n}{2}}\ll q(\eta)^{\frac{n}{2}+\epsilon},$
where $\omega(a)$ is the number of distinct primes of $a$.
Case 2: $n$ odd: We use (11.11) and (11.12) and show by case by case.
Now recall the notations from Section 4.2: Since $V$ has signature $(2,n)$, we can write $q(x,y)=q_{2}(x)-q_{n}(y)$, where $V^{\prime}=V_{2}\oplus V_{n}$. Recall that $\{v_{1},v_{2}\}$ is an orthonormal basis of $V_{2}(\mathbb{R})$. Now write $\eta=a_{1}v_{1}+a_{2}v_{2}+w$, where $w\in V_{n}$. Then $q(\eta)=a_{1}^{2}+a_{2}^{2}-q_{n}(w)$, and $|\langle\eta,v_{1}+iv_{2}\rangle|^{2}=a_{1}^{2}+a_{2}^{2}$. Therefore, the number of $\eta\in V^{\prime}(\mathbb{Z})$ such that $|\langle\eta,v_{1}+iv_{2}\rangle|=T$ and $q(\eta)>0$, is $O(T^{\frac{n}{2}})$. Now $K_{v}(z)\sim\sqrt{\frac{\pi}{2}}\frac{e^{-z}}{\sqrt{z}}$ as $|z|\to\infty$. Hence the formal series $F_{f}(g)$ is absolutely convergent and also is of moderate growth for any fixed norm $g_{\infty}||$ of $g_{\infty}\in G(\mathbb{R})$. ∎
Lemma 8.3.
The formal series $F_{f}(g)=F(g;\{\alpha_{p}\}_{p})$ is left $P(\mathbb{Q})$-invariant.
Proof.
Let $\gamma={\rm diag}(c,m,c^{-1})n\in P(\mathbb{Q})=M(\mathbb{Q})N(\mathbb{Q})$. By (8.1) and (3.2), we have
$F_{f}(\gamma g)=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}A(\eta;\{\alpha_{p}\}_{p})(\gamma g_{\mathbb{f}})\mathcal{W}_{2\pi\eta}(\gamma g_{\infty})$
$\phantom{xxxxx}=\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}|c|^{l-\frac{n}{2}}_{\mathbb{A}_{f}}\cdot|c|^{\frac{n+2}{2}}_{\mathbb{A}_{f}}A(\gamma^{-1}\eta;\{\alpha_{p}\}_{p})(g_{\mathbb{f}})|c|^{(l+1)}_{\infty}\mathcal{W}_{2\pi\gamma^{-1}\eta}(g_{\infty}).$
Clearly, $|c|^{l-\frac{n}{2}}_{\mathbb{A}_{f}}\cdot|c|^{\frac{n+2}{2}}_{\mathbb{A}_{f}}|c|^{(l+1)}_{\infty}=|c|^{l+1}_{\mathbb{A}}=1$ since $c\in\mathbb{Q}^{\times}$. Replacing $\gamma^{-1}\eta$ with $\eta$ does not change the conditions in the summation. Thus, we have the claim. ∎
Proposition 8.4.
The formal series $F_{f}(g)=F(g;\{\alpha_{p}\}_{p})$ is left $H^{{\rm ss}}(\mathbb{Q})$-invariant.
Proof.
[…]
where $\Phi$ runs over $\mathcal{S}(X(\mathbb{A}))$. Notice $H^{{\rm ss}}(\mathbb{Q})$ splits under the double covering $\widetilde{H^{{\rm ss}}(\mathbb{A})}\longrightarrow H^{{\rm ss}}(\mathbb{A})$ and $\Theta_{\psi_{S}}(\widetilde{h}v)$ is left $H^{{\rm ss}}(\mathbb{Q})$-invariant. Then, by the argument for [16, Lemma 5.4], for any fixed $v$ ($l\leq v\leq l$), we may check the $H^{{\rm ss}}(\mathbb{Q})$-invariance of the $v$-th component $((F_{f})_{S,\Phi})_{v}$ (defined as (7.19)) of $(F_{f})_{S,\Phi}$ for any $\Phi\in\mathcal{S}(X(\mathbb{A}))$.
Now we choose $\Phi_{\infty}$ as in Section 8.2 for each $v\in\{-l,0,l\}$. By a similar computation proving (7.16) and Corollary 7.5, we have
$((F_{f})_{S,\Phi})_{v}=\sum_{\xi\in X(\mathbb{Q})}((F_{f})_{\xi,S,\Phi})_{v}$
where if $S>0$ and $v\in\{-l,0\}$,
$(F_{f})_{\xi,S,\Phi}(\widetilde{h})_{v}=C_{4}(S,\Phi_{\infty})\varepsilon^{n}_{\infty}\sqrt{\beta}^{l-\frac{n}{2}+1}\Bigg(\sum_{a\in\mathbb{Q}_{>0}\atop a+S\sigma(\xi,\xi)\geq 0}A(\eta_{a,S,\xi};\{\alpha_{p}\}_{p})(h_{\mathbb{f}})(\bar{q}_{\tau})^{a+S\sigma(\xi,\xi)}\Bigg)$
and if $S<0$ and $v\in\{0,l\}$,
$(F_{f})_{\xi,S,\Phi}(\widetilde{h})_{v}=C_{4}(S,\Phi_{\infty})\varepsilon^{n}_{\infty}\sqrt{\beta}^{l-\frac{n}{2}+1}\Bigg(\sum_{a\in\mathbb{Q}_{<0}\atop a+S\sigma(\xi,\xi)<0}A(\eta_{a,S,\xi};\{\alpha_{p}\}_{p})(h_{\mathbb{f}})(q_{\tau})^{-(a+S\sigma(\xi,\xi))}\Bigg).$
Since $F_{f}$ is right G(^)G($\widehat{\mathbb{Z}}$) by construction and left $P(\mathbb{Q})$-invariant by Lemma 8.3, we may assume $\Phi_{\mathbb{f}}={\rm char}_{V^{\prime}(\widehat{\mathbb{Z}})}$. Assume $S<0$. First, we observe the $l$-th component $((F_{f})_{\xi,S,\Phi})_{l}$ for $\Phi$ whose archimedean component $\Phi_{\infty}$ is Gaussian $\Phi_{S,0}$. Then, Applying Section 8.2 and [11, Theorem 10.2] for $n$ even and [10, Theorem 10.5] for $n$ odd, the $l$-th component $((F_{f})_{\xi,S,\Phi})_{l}$ is automorphic, hence left $H^{\rm ss}(\mathbb{Q})$-invariant. It is also corresponding to a holomorphic modular form of weight $l-\frac{n}{2}+1$. Further, $C_{4}(S,\Phi_{\infty})=C_{2}(S,\Phi_{S,0})$ is non-zero by Remark 7.1. Thus, the formal Fourier-Jacobi coefficient
(8.4) ${\rm FJC}_{S,\xi}(\widetilde{h}):=\sqrt{\beta}^{l-\frac{n}{2}+1}\Bigg(\sum_{a\in\mathbb{Q}_{<0}\atop a+S\sigma(\xi,\xi)<0}A(\eta_{a,S,\xi};\{\alpha_{p}\}_{p})(h_{\mathbb{f}})(q_{\tau})^{-(a+S\sigma(\xi,\xi))}\Bigg)$
corresponds to a holomorphic modular form of weight $l-\frac{n}{2}+1$ in the sense of Remark 7.8. Notice that it is independent of $\Phi_{\infty}$.
Now we turn to look the 0-th component $((F_{f})_{\xi,S,\Phi})_{0}$ for any $\Phi\in\mathcal{S}(X(\mathbb{A}))$ with $\Phi_{\mathbb{f}}={\rm char}_{V^{\prime}(\widehat{\mathbb{Z}})}$. By the above computation we have
[…]
Theorem 8.5.
Keep the notation being as above. The formal series $F_{f}$ is automorphic cuspidal form on $G(\mathbb{A})$ of weight $l+1$ with respect to G(^)G($\widehat{\mathbb{Z}}$) in the sense of Section 3.
Proof.
Since $G(\mathbb{Q})=H^{{\rm ss}}(\mathbb{Q})P(\mathbb{Q})$, the claim follows from Lemma 8.2 and Lemma 8.3, and Proposition 8.4. ∎
8.4. Hecke Eigenform Property
[…]
Lemma 8.6.
The Eisenstein series $E_{l}$ is a Hecke eigenform.
Proof.
We may prove it for $T=T_{h}\in\mathcal{H}_{\mathbb{f}}$. By definition, we have
[…]
Since the finite part of $f$ is a non-zero section of the one-dimensional space $\Big({\rm Ind}^{G(\mathbb{A}_{\mathbb{f}})}_{P(\mathbb{A}_{\mathbb{f}})}|\nu|_{\mathbb{f}}^{l-\frac{n+2}{2}}\Big)^{G(\widehat{\mathbb{Z}})}$ which also preserved under the action of $\mathcal{H}_{\mathbb{f}}$. Thus, there exists $c_{h}\in\mathbb{C}^{\times}$ such that $T_{h}f=c_{h}f$. The claim follows. ∎
Proposition 8.7.
The automorphic form $F_{f}$ is a Hecke eigenform.
Proof.
We may prove it for $T=T_{h},\ h\in P(\mathbb{Q})$ . We write $G(\widehat{\mathbb{Z}})hG(\widehat{\mathbb{Z}})=\displaystyle\coprod_{i=1}^{r}h_{i}G(\widehat{\mathbb{Z}}),\ h_{i}\in P(\mathbb{Q})$. It is easy to see that $|\nu(h)|_{\mathbb{f}}=|\nu(h_{i})|_{\mathbb{f}}$ for any $i$. Let $l^{\prime}>n+1$ be a sufficiently large integer. Assume $n$ is even. Then, we see that
$T_{h}E_{l^{\prime}}(g)=\sum_{i}E_{l^{\prime}}(g_{\infty},g_{\mathbb{f}}h_{i}\Phi_{\mathbb{f}},s=l^{\prime}+1)=C_{l^{\prime},n}\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}\sum_{i=1}^{r}q(\eta)^{\frac{l^{\prime}-\frac{n}{2}}{2}}\prod_{p<\infty}\widetilde{Q}_{\eta,(g_{\mathbb{f}}h_{i})\cdot\Phi_{p}}(p^{\frac{l^{\prime}-\frac{n}{2}}{2}})W_{2\pi\eta}(g_{\infty})$
where $\widetilde{Q}_{\eta,\Phi_{p}}=\widetilde{Q}_{\eta,p}$. By Lemma 4.1, we have
$T_{h}E_{l^{\prime}}(g)=C_{l^{\prime},n}|\nu(h)|_{\mathbb{f}}^{-\frac{n+2}{2}}\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}\sum_{i=1}^{r}q(\eta)^{\frac{l^{\prime}-\frac{n}{2}}{2}}\prod_{p<\infty}\widetilde{Q}_{h^{-1}_{i}\eta,g_{\mathbb{f}}\cdot\Phi_{p}}(p^{\frac{l^{\prime}-\frac{n}{2}}{2}})W_{2\pi\eta}(g_{\infty}).$
We choose $\eta_{0}\in V^{\prime}(\mathbb{Q})$ in Lemma 8.1. By Lemma 8.6, there exists a non-zero constant $c_{h}$ such that $T_{h}E_{l^{\prime}}(g)=c_{h}E_{l^{\prime}}(g)$. Evaluating $g=g_{\infty}$ (hence $g_{\mathbb{f}}=1$), we see that $c_{h}=\alpha_{h}(\{p^{\frac{l^{\prime}-\frac{n}{2}}{2}}\}_{p})$ where
$\alpha_{h}(X):=|\nu(h)|_{\mathbb{f}}^{-\frac{n+2}{2}}\sum_{i=1}^{r}\prod_{p<\infty}\widetilde{Q}_{h^{-1}_{i}\eta_{0},\Phi_{p}}(X_{p})\in\otimes^{\prime}_{p<\infty}\mathbb{C}[X_{p},X^{-1}_{p}],\ X=\{X_{p}\}_{p}.$
By using [10, Lemma 10.1] and the argument for [15, Proposition 10.1], we see that
$T_{h}F_{f}=\alpha_{h}(\{\alpha_{p}\}_{p})F_{f}.$
The odd case is similarly proved and we omit the proof. ∎
9. $L$-functions of Ikeda type lift
Let $f=\displaystyle\sum_{m=1}^{\infty}a_{f}(m)m^{\frac{k-1}{2}}q^{m}$ be a Hecke eigen cusp form of weight $k=\left\{\begin{array}[]{ll}l-\frac{n}{2}+1&\text{for $n$ even,}\\ 2l-n+1&\text{for $n$ odd}\end{array}\right$. with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $L(s,\pi_{f})=\displaystyle\prod_{p}L(s,\pi_{p})$, where $L(s,\pi_{p})=(1-\alpha_{p}p^{-s})^{-1}(1-\alpha_{p}^{-1}p^{-s})^{-1}$.
9.1. $n$ even
Let $n=8a+2$ ($n>2$). Let $F=F_{f}$ be the Ikeda type lift of $f$ on ${\rm SO}(3,n+1)$, and let $\Pi_{F}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F$.
When $p$ is unramified, $G(\mathbb{Q}_{p})\simeq{\rm SO}(4a+3,4a+3)$, and $\Pi_{p}={\rm Ind}_{P}^{G}\,|\ |^{2s_{p}}\rtimes 1_{{\rm SO}}$, where $p^{s_{p}}=\alpha_{p}$ and $1_{{\rm SO}}$ is the trivial representation of ${\rm SO}(4a+2,4a+2)$, i.e., the quotient of ${\rm Ind}_{B}^{{\rm SO}(4a+2,4a+2)}\,|\ |^{4a+1}\otimes|\ |^{4a}\otimes\cdots\otimes|\ |^{0}$, where $B$ is the Borel subgroup of ${\rm SO}(4a+2,4a+2)$. Notice $2s_{p}$ since we replace $p^{\frac{l-\frac{n}{2}}{2}}$ by $\alpha_{p}$. If $n$ is odd, we replace $p^{l-\frac{n}{2}}$ by $\alpha_{p}$.
Consider $\widetilde{G}={\rm SO}(4a+4,4a+4)$, and $R=M^{\prime}N^{\prime}$, $M^{\prime}\simeq{\rm GL}_{1}\times{\rm SO}(4a+3,4a+3)$. Then Langlands-Shahidi theory applies to $(\widetilde{G},M^{\prime},\Pi_{p})$, and we have
Theorem 9.1.
The standard $L$-function of $\Pi_{F}$ is given by
$L(s,\Pi_{F})=L(s,{\rm Sym}^{2}\pi_{f})\zeta(s+\frac{n}{2})\zeta(s+\frac{n}{2}-1)\cdots\zeta(s+1)\zeta(s)\zeta(s-1)\cdots\zeta(s-\frac{n}{2}).$
Remark 9.2.
In [19] Y. Li, N. Narita and A. Pitale constructed non-tempered cusp forms on $O(1,8n+1)$ from Maass cusp forms. Their $L$-function has the exactly same shape.
Remark 9.3.
We give a conjectural Arthur parameter of $\Pi_{F}$: Let $\phi_{f}:\mathcal{L}\longrightarrow{\rm SL}_{2}(\mathbb{C})$ be the hypothetical Langlands parameter attached to $f$. Then ${\rm Sym}^{2}\phi_{f}:SL_{2}(\mathbb{C})\times{\rm SO}(3,\mathbb{C})$ parametrizes ${\rm Sym}^{2}\pi_{f}$. The distinguished unipotent orbit $(8a+3)$ of ${\rm SO}(8a+3,\mathbb{C})$ gives rise to a map $\phi_{u}:{\rm SL}_{2}(\mathbb{C})\longrightarrow{\rm SO}(8a+3,\mathbb{C})$. Then consider
$\phi={\rm Sym}^{2}\phi_{f}\oplus\phi_{u}:\mathcal{L}\times{\rm SL}_{2}(\mathbb{C})\longrightarrow{\rm SO}(3,\mathbb{C})\times{\rm SO}(8a+3,\mathbb{C})\subset{\rm SO}(8a+6,\mathbb{C})={}^{L}G.$
We expect that $\phi$ parametrizes $\Pi_{F}$. The global A-parameter of $\Pi_{F}$ [28] is
${\rm Sym}^{2}\pi_{f}[1]\oplus 1[8a+3].$
9.2. $n$ odd
Let $n=8a+3$. Let $F=F_{f}$ be the Ikeda type lift of $f$ on ${\rm SO}(3,1+n)$, and let $\Pi_{F}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F$.
When $p$ is unramified, $G(\mathbb{Q}_{p})\simeq{\rm SO}(4a+3,4a+4)$, and $\Pi_{p}={\rm Ind}_{P}^{G}\,|\ |^{s_{p}}\rtimes 1_{{\rm SO}}$, where $p^{s_{p}}=\alpha_{p}$ and $1_{{\rm SO}}$ is the trivial representation of ${\rm SO}(4a+2,4a+3)$, i.e., the quotient of ${\rm Ind}_{B}^{{\rm SO}(4a+2,4a+3)}\,|\ |^{4a+\frac{3}{2}}\otimes|\ |^{4a+\frac{1}{2}}\otimes\cdots\otimes|\ |^{\frac{1}{2}}.$
Consider $\widetilde{G}={\rm SO}(4a+4,4a+5)$, and $R=M^{\prime}N^{\prime}$, $M^{\prime}\simeq{\rm GL}_{1}\times{\rm SO}(4a+3,4a+4)$. Then Langlands-Shahidi theory applies to $(\widetilde{G},M^{\prime},\Pi_{p})$, and we have
Theorem 9.4.
The standard $L$-function of $\Pi_{F}$ is given by
$L(s,\Pi_{F})=L(s,\pi_{f})\zeta(s+\frac{n}{2})\zeta(s+\frac{n}{2}-1)\cdots\zeta(s+\frac{1}{2})\zeta(s-\frac{1}{2})\cdots\zeta(s-\frac{n}{2}).$
If $n=8a+3$, $^{L}G={\rm Sp}(8a+6,\mathbb{C})$. The distinguished unipotent orbit $(8a+4)$ of ${\rm Sp}(8a+4,\mathbb{C})$ gives rise to a map $\phi_{u}:{\rm SL}_{2}(\mathbb{C})\longrightarrow{\rm Sp}(8a+4,\mathbb{C})$. Then we expect that
$\phi=\phi_{f}\oplus\phi_{u}:\mathcal{L}\times{\rm SL}_{2}(\mathbb{C})\longrightarrow{\rm SL}_{2}(\mathbb{C})\times{\rm Sp}(8a+4,\mathbb{C})\subset{\rm Sp}(8a+6,\mathbb{C}),$
parametrizes $\Pi_{F}$. The global A-parameter of $\Pi_{F}$ [28] is
$\pi_{f}[1]\oplus 1[8a+4].$
10. Ikeda type lift as CAP representation
Let $\Pi_{F}=\otimes_{p}^{\prime}\Pi_{p}$ be the Ikeda type lift in the previous section. Then $\Pi_{F}$ is not a CAP representation in the usual sense since there are not many $\mathbb{Q}$-parabolic subgroups of ${\rm SO}(3,n+1)$. We show that $\pi_{F}$ is a CAP representation in a more general sense, namely, there exists a parabolic group $R=MN$ of split $G={\rm SO}(\frac{n+4}{2},\frac{n+4}{2})$ or ${\rm SO}(\frac{n+3}{2},\frac{n+5}{2})$, and a cuspidal representation $\tau=\otimes_{p}\tau_{p}$ of $M(\mathbb{A}_{\mathbb{Q}})$ and a parameter $\Lambda_{0}$ such that for all finite prime p, $\Pi_{p}$ is a quotient of ${\rm Ind}_{R(\mathbb{Q}_{p})}^{G(\mathbb{Q}_{p})}\tau_{p}\otimes\exp(\Lambda_{0},H_{R}(\,))$.
Case: $n$ even. Let $n=8a+2$. Let $R=MN$, and $M\simeq\underbrace{{\rm GL}_{1}\times\cdots\times{\rm GL}_{1}}_{4a+1}\times{\rm SO}(2,2)$, and we use the fact that there is a homomorphism ${\rm SL}_{2}\times{\rm SL}_{2}\longrightarrow{\rm SO}(2,2)$ whose kernel is isomorphic to $\mathbb{Z}/2\mathbb{Z}$. In particular, $(a,a^{-1})\times(b,b^{-1})$ maps to $(ab,ab^{-1},a^{-1}b,a^{-1}b^{-1})$. Hence $\sigma=|\det|^{-\frac{1}{2}}\pi_{f}\otimes|\det|^{\frac{1}{2}}\pi_{f}=\otimes_{p}^{\prime}\sigma_{p}$ can be considered as a cuspidal representation of ${\rm SO}(2,2)$. Under the homomorphism, the Satake parameter ${\rm diag}(p^{-\frac{1}{2}}\alpha_{p},p^{-\frac{1}{2}}\alpha_{p}^{-1})\times{\rm diag}(p^{\frac{1}{2}}\alpha_{p},p^{\frac{1}{2}}\alpha_{p}^{-1})$ of ${\rm GL}_{2}\times{\rm GL}_{2}$ corresponds to ${\rm diag}(1,\alpha_{p}^{2},\alpha_{p}^{-2},1)$ in ${\rm SO}(2,2)$. Then $\Pi_{p}$ is a quotient of ${\rm Ind}_{R}^{G}\,|\ |^{4a+1}\otimes\cdots\otimes|\ |\otimes\sigma_{p}$, which is equivalent to ${\rm Ind}_{R}^{G}\,|\ |^{4a+1}\otimes\cdots\otimes|\ |\otimes|\ |^{0}\otimes|\ |^{2s_{p}}$, where $s_{p}$ is as in Section 9.
Remark 10.1.
When $n=2$, then $G={\rm SO}(3,3)$ and our method (in Section 6) breaks down, and we do not get a cusp form on ${\rm SO}(3,3)$. For example, it is no longer true that ${\rm SO}(3,3)$ has a relative root system of type $B_{3}$ in Section 6.1. Thus, we cannot apply the results in Section 8.2. It agrees with the fact that ${\rm SL}_{4}$ is isogenous to ${\rm SO}(3,3)$, and ${\rm SL}_{4}$ does not have CAP representations, since any cuspidal representation of ${\rm SL}_{n}$ can be extended to a cuspidal representation of ${\rm GL}_{n}$ (cf. [3, p.236, line 9-10]). On the other hand, H. Narita, A. Pitale and S. Wagh [21] constructed a CAP representation on ${\rm O}(1,5)$ from Maass cusp form using the theta lift.
Case: $n$ odd. Let $n=8a+3$. Let $R=MN$, and $M\simeq\underbrace{{\rm GL}_{1}\times\cdots\times{\rm GL}_{1}}_{4a+1}\times{\rm SO}(2,1)$, and we use the fact that ${\rm SO}(2,1)\simeq{\rm PGL}_{2}$. Hence $\pi_{f}=\otimes_{p}^{\prime}\pi_{p}$ can be considered as a representation of ${\rm SO}(2,1)$. Then $\Pi_{p}$ is a quotient of ${\rm Ind}_{R}^{G}\,|\ |^{4a+\frac{3}{2}}\otimes\cdots\otimes|\ |^{\frac{1}{2}}\otimes\pi_{p}$.
11. Siegel series
In this section we compute the Siegel series from Section 4.1.
11.1. The case when $n$ is even
Now suppose $\dim\,V^{\prime}=n+2=2m$. Since our orthogonal group splits at any $p$, we may assume that
$q(x_{1},...,x_{m},y_{1},...,y_{m})=x_{1}y_{1}+\cdots+x_{m}y_{m}.$
Let $\eta=(a_{1},...,a_{m},b_{1},...,b_{m})$. Then the above sum is
(11.1) $B_{r,\eta}=\sum_{u_{1},...,u_{m},v_{1},...,v_{m}\in\mathbb{Z}/p^{r}\mathbb{Z}\atop u_{1}v_{1}+\cdots+u_{m}v_{m}\equiv 0\,\text{mod $p^{r}$}}e^{2\pi i\frac{1}{p}(a_{1}v_{1}+\cdots+a_{m}v_{m}+b_{1}u_{1}+\cdots+b_{m}u_{m})}.$
Now note that
$\sum_{x\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{2\pi i\frac{x}{p^{r}}(u_{1}v_{1}+\cdots+u_{m}v_{m})}=\begin{cases}p^{r},&\text{if $u_{1}v_{1}+\cdots+u_{m}v_{m}\equiv 0$ (mod $p^{r}$)}\\ 0,&\text{otherwise.}\end{cases}$
Hence the above sum is
(11.2) $p^{-r}\sum_{u_{1},...,u_{m},v_{1},...,v_{m}\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{2\pi i\frac{1}{p^{r}}(a_{1}v_{1}+\cdots+a_{m}v_{m}+b_{1}u_{1}+\cdots+b_{m}u_{m})}\sum_{x\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{2\pi i\frac{x}{p^{r}}(u_{1}v_{1}+\cdots+u_{m}v_{m})}$
(11.3) $\phantom{xxxxxx}=p^{-r}\sum_{x\in\mathbb{Z}/p^{r}\mathbb{Z}}\prod_{i=1}^{m}\left(\sum_{u_{i},v_{i}\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{2\pi i\frac{x}{p^{r}}(u_{i}v_{i})}e^{2\pi i\frac{1}{p^{r}}(a_{i}v_{i}+b_{i}u_{i})}\right).$
Now we have
Lemma 11.1.
Let $j=v_{p}(x)$. Then
$\sum_{u,v\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{2\pi i\frac{1}{p^{r}}(xuv+av+bu)}=\begin{cases}p^{2r},&\text{if $j\geq r$, $p^{r}|a$, $p^{r}|b$}\\ p^{r+j}e^{2\pi i\frac{-a^{\prime}b^{\prime}x_{0}^{-1}}{p^{r-j}}},&\text{if $0\leq j<r$, $p^{j}|a$, $p^{j}|b$}\\ 0,&\text{otherwise}\end{cases},$
where $x=p^{j}x_{0}$ with $p\nmid x_{0}$ and $x_{0}^{-1}$ is taken modulo $p^{r-j}$, and $a=p^{j}a^{\prime}$, $b=p^{j}b^{\prime}$ when $j<r$.
Proof.
Let $f(x,a,b)$ be the exponential sum. Put $N=p^{r}$, and $\zeta_{N}=e^{2\pi i/N}$. Then
[…]
Case A: $j\geq r$. Then if $p^{r}\nmid a$, it is clear that $f(x,a,b)=0$. Suppose $p^{r}|a$. Then it is easy to see
$f(x,a,b)=N\sum_{\text{$u$ (mod $N$)}}\zeta_{N}^{bu}=\begin{cases}N^{2}=p^{2r},&\text{if $p^{r}|b$}\\ 0,&\text{otherwise}\end{cases}.$
In conclusion, we have, if $j\geq r$,
$f(x,a,b)=\begin{cases}p^{2r},&\text{if $p^{r}|a$ and $p^{r}|b$}\\ 0,&\text{otherwise}\end{cases}.$
Case B: $0\leq j<r$. The congruence $xu\equiv-a$ (mod $N$) is solvable if and only if $p^{j}|a$. Then the solutions are
$\text{$u\equiv u_{0}$ (mod $p^{r-j}$), $u_{0}\equiv-a^{\prime}x_{0}^{-1}$ (mod $p^{r-j}$)},$
and there are exactly $p^{j}$ of them: $u=u_{0}+tp^{r-j}$ with $t=0,1,...,p^{j}-1$. Hence
$f(x,a,b)=N\sum_{t=0}^{p^{j}-1}\zeta_{N}^{b(u_{0}+tp^{r-j})}=N\zeta_{N}^{bu_{0}}\sum_{t=0}^{p^{j}-1}e^{2\pi i\frac{bt}{p^{j}}}.$
Therefore,
$f(x,a,b)=\begin{cases}Np^{j}\zeta_{N}^{bu_{0}}=p^{r+j}e^{2\pi i\frac{b^{\prime}u_{0}}{p^{r-j}}},&\text{if $p^{j}|a$ and $p^{j}|b$}\\ 0,&\text{otherwise}\end{cases}.$
This proves the lemma. ∎
So if $j<r$, the sum over elements $x=p^{j}c$, $p\nmid c$ in (11.3) is
[…]
By setting $a_{i}=p^{j}a_{i}^{\prime},b_{i}=p^{j}b_{i}^{\prime}$ for each $i$, the above sum is,
(11.4) $\sum_{c\in\mathbb{Z}/p^{r-j}\mathbb{Z},\,p\nmid c}e^{2\pi i\frac{1}{p^{r-j}}c(a_{1}^{\prime}b_{1}^{\prime}+\cdots+a_{m}^{\prime}b_{m}^{\prime})}.$
It is the Ramanujan sum. By [13, p. 44], it is
$\sum_{d|(p^{r-j},A_{j})}d\mu\Big(\frac{p^{r-j}}{d}\Big),$
where $\mu(n)$ is the Möbius function and $A_{j}=a_{1}^{\prime}b_{1}^{\prime}+\cdots+a_{m}^{\prime}b_{m}^{\prime}$. Then
[…]
Therefore, if $\eta\equiv 0$ (mod $p^{r}$), then (11.3) is
$p^{2rm-r}+\sum_{j=0}^{r-1}p^{rm-r+jm}(p^{r-j}-p^{r-j-1})=p^{2rm-r}+\sum_{j=0}^{r-1}(p^{rm+jm-j}-p^{rm+jm-j-1}).$
Hence, $J(s,\eta,\Phi_{p})=0$ unless $\eta\in V^{\prime}(\mathbb{Z}_{p})$. Now let $p^{k}\|\eta$, and $p^{2k+k^{\prime}}\|q(\eta)$ for some $k^{\prime}$. Then
$J(s,\eta,\Phi_{p})=\sum_{r=0}^{\infty}p^{-rs}B_{r,\eta},$
where $B_{0,\eta}=1$, and if $r\leq k$, since $\eta\equiv 0$ (mod $p^{r}$),
$B_{r,\eta}=p^{2rm-r}+\sum_{j=0}^{r-1}(p^{rm+jm-j}-p^{rm+jm-j-1}).$
If $r>k$, (11.1) is
$B_{r,\eta}=\sum_{j=0}^{k}p^{rm-r+jm}\sum_{c\in\mathbb{Z}/p^{r-j}\mathbb{Z},\,p\nmid c}e^{2\pi i\frac{1}{p^{r+j}}2c(a_{1}b_{1}+\cdots+a_{m}b_{m})}=\sum_{j=0}^{k}p^{rm-r+jm}\left(\sum_{d|(p^{r-j},A_{j})}d\mu\Big(\frac{p^{r-j}}{d}\Big)\right),$
where $A_{j}=p^{-2j}q(\eta)$.
If $r\geq 2k+k^{\prime}+2$, $B_{r,\eta}=0$. And if $r=2k+k^{\prime}+1$,
$B_{2k+k^{\prime}+1,\eta}=-p^{(2k+k^{\prime}+1)m-1}.$
If $r=2k+k^{\prime}$,
$B_{2k+k^{\prime},\eta}=\begin{cases}p^{k^{\prime}m-k^{\prime}}(p^{k^{\prime}}-p^{k^{\prime}-1}),&\text{if $k=0$}\\ p^{rm-r}(p^{2k+k^{\prime}}-p^{2k+k^{\prime}-1})+p^{rm-r+m}(-p^{2k+k^{\prime}-2})=p^{rm}-p^{rm-k^{\prime}}-p^{rm+m-2},&\text{if $k>0$.}\end{cases}$
If $k<r\leq k+k^{\prime}$, then $r-j\leq k+k^{\prime}-j\leq 2k+k^{\prime}-2j$. Hence
$\sum_{d|(p^{r-j},A_{j})}d\mu\Big(\frac{p^{r-j}}{d}\Big)=p^{r-j}-p^{r-j-1}.$
Therefore,
$B_{r,\eta}=\sum_{j=0}^{k}p^{rm-r+jm}(p^{r-j}-p^{r-j-1}).$
If $k+k^{\prime}<r\leq 2k+k^{\prime}-1$, then $r-j\leq 2k+k^{\prime}-2j$ implies $j\leq 2k+k^{\prime}-r$. So
$B_{r,\eta}=\sum_{j=0}^{2k+k^{\prime}-r}p^{rm-r+jm}(p^{r-j}-p^{r-j-1})+p^{rm-r+m(2k+k^{\prime}-r+1)}(-p^{2r-2k-k^{\prime}-2})$
$=\sum_{j=0}^{2k+k^{\prime}-r}p^{rm}(p^{j(m-1)}-p^{j(m-1)-1})-p^{r+(2k+k^{\prime})(m-1)+m-2}.$
This is valid also for $r=2k+k^{\prime}$.
When $r=1$,
[…]
If $\eta\equiv 0$ (mod $p$), $B_{1,\eta}=C(n)=\#\{u\in V^{\prime}(\mathbb{Z}/p\mathbb{Z})|\,q(u)\equiv 0\,\text{mod $p$}\}$. This gives another proof of Lemma 4.4.2 and Lemma 4.4.3 in [23].
Proof of Theorem 4.2. Consider $J(s,\eta,\Phi_{p})(1-p^{m-1-s})^{-1}$. Namely,
$\left(\sum_{r=0}^{\infty}B_{r,\eta}p^{-rs}\right)(1-p^{m-1-s})^{-1}=\sum_{a=0}^{\infty}\left(\sum_{r=0}^{a}B_{r,\eta}p^{(m-1)(a-r)}\right)p^{-as}.$
We claim that if $a\geq 2k+k^{\prime}+1$, $\sum_{r=0}^{a}B_{r,\eta}p^{(m-1)(a-r)}=0$. Then this shows that
$Q_{\eta,p}(p^{m-s})=\sum_{a=0}^{2k+k^{\prime}}\left(\sum_{r=0}^{a}B_{r,\eta}p^{(m-1)(a-r)}\right)p^{-as}.$
First, consider the case $k=0$. Then if $a\geq k^{\prime}+1$,
$\sum_{r=0}^{a}B_{r,\eta}p^{-r(m-1)}=1+\sum_{r=1}^{k^{\prime}-1}(p^{r}-p^{r-1})+p^{k^{\prime}}(1-p^{-1})-p^{k^{\prime}}=0.$
Now let $k>0$. If $a\geq 2k+k^{\prime}+1$,
$\sum_{r=0}^{a}B_{r,\eta}p^{-r(m-1)}=\sum_{r=0}^{k}\left(p^{rm}+\sum_{j=0}^{r-1}(p^{j(m-1)+r}-p^{j(m-1)+r-1})\right)+\sum_{r=k+1}^{k+k^{\prime}}\sum_{j=0}^{k}(p^{j(m-1)+r}-p^{j(m-1)+r-1})$
$+\sum_{r=k+k^{\prime}+1}^{2k+k^{\prime}}\left(\sum_{j=0}^{2k+k^{\prime}-r}(p^{j(m-1)+r}-p^{j(m-1)+r-1})-p^{-rm+2r+(2k+k^{\prime})(m-1)+m-2}\right)-p^{2k+k^{\prime}}.$
By Mathematica, we can show that it is zero. By Mathematica, we can show that $Q_{\eta,p}(p^{m-s})$ is a polynomial in $p^{m-s}$:
$Q_{\eta,p}(p^{m-s})=\sum_{a=0}^{2k+k^{\prime}}A_{a,\eta}(p^{m-s})^{a},$
where
(11.5) $A_{a,\eta}=\begin{cases}1+p^{m-1}+\cdots+(p^{m-1})^{a},&\text{if $0\leq a\leq k$,}\\ 1+p^{m-1}+\cdots+(p^{m-1})^{k},&\text{if $k+1\leq a\leq k+k^{\prime}$,}\\ 1+p^{m-1}+\cdots+(p^{m-1})^{2k+k^{\prime}-a},&\text{if $k+k^{\prime}+1\leq a\leq 2k+k^{\prime}$.}\end{cases}$
Let $X=p^{m-s}$, and we can write
$Q_{\eta,p}(X)=\sum_{a=0}^{k-1}A_{a,\eta}(X^{a}+X^{2k+k^{\prime}-a})+\frac{1}{2}A_{k,\eta}\sum_{a=k}^{k+k^{\prime}}(X^{a}+X^{2k+k^{\prime}-a}).$
Then it satisfies the functional equation $X^{2k+k^{\prime}}Q_{\eta,p}(X^{-1})=Q_{\eta,p}(X).$
11.2. The case when $n$ is odd
If $n$ is odd, let ${\rm dim}\,V^{\prime}=n+2=2m+1$. In this case, we need to modify the previous argument. Since the orthogonal group splits at $p$, we may assume that
$q(u_{1},...,u_{m},v_{1},...,v_{m},u_{0})=u_{1}v_{1}+\cdots+u_{m}v_{m}+u_{0}^{2}.$
For $\eta=(a_{1},...,a_{m},b_{1},...,b_{m},a_{0})$, $u=(u_{1},...,u_{m},v_{1},...,v_{m},u_{0})$,
$(\eta,u)=a_{1}v_{1}+\cdots+a_{m}v_{m}+b_{1}u_{1}+\cdots+b_{m}u_{m}+2a_{0}u_{0}.$
Then
$B_{r,\eta}=p^{-r}\sum_{x\in\mathbb{Z}/p^{r}\mathbb{Z}}\left(\sum_{u_{0}\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{\frac{2\pi i}{p^{r}}(xu_{0}^{2}+2a_{0}u_{0})}\right)\prod_{i=1}^{m}\left(\sum_{u_{i},v_{i}\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{2\pi i\frac{x}{p^{r}}(u_{i}v_{i})}e^{2\pi i\frac{1}{p^{r}}(a_{i}v_{i}+b_{i}u_{i})}\right).$
Now we need to divide into two cases.
11.2.1. The case when $p>2$ (and $n$ is odd)
Recall the following exponential sums [27], a generalization of the Ramanujan sum (11.4): Let $p$ be an odd prime. Let $\sum_{\text{$n$ mod $p^{\alpha}$}}^{\prime}$ mean the sum $\sum_{\text{$n$ mod $p^{\alpha}$}\atop p\nmid n}$. Then if $\alpha$ is odd, and we put $h_{1}=\frac{h}{p^{\alpha-1}}$ when $p^{\alpha-1}\|h$, we have
[…]
Now for $x=p^{j}c$ with $p\nmid c$ and $j<r$. Then $\sum_{u_{0}\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{\frac{2\pi i}{p^{r}}(xu_{0}^{2}+a_{0}u_{0})}=0$ unless $a_{0}\equiv 0$ (mod $p^{j}$). Let $a_{0}=p^{j}a_{0}^{\prime}$. Let $c^{-1}$ be taken modulo $p^{r-j}$. Then
(11.6) $\sum_{u_{0}\in\mathbb{Z}/p^{r}\mathbb{Z}}e^{\frac{2\pi i}{p^{r}}(xu_{0}^{2}+2a_{0}u_{0})}=e^{-2\pi i\frac{1}{p^{r-j}}(a_{0}^{\prime})^{2}c^{-1}}p^{\frac{r+j}{2}}\begin{cases}1,&\text{if $r-j$ even}\\ (\frac{c}{p})\sqrt{(\frac{-1}{p})},&\text{if $r-j$ odd}\end{cases}.$
In summary, we have the following: Recall
$J(s,\eta,\Phi_{p})=\sum_{r=0}^{\infty}p^{-rs}B_{r,\eta},\quad B_{0,\eta}=1.$
Suppose $p^{k}\|\eta$ and $p^{2k+k^{\prime}}\|q(\eta)$. Let $q(\eta)_{1}=q(\eta)/p^{2k+k^{\prime}}$. If $k\geq r$, then $\eta\equiv 0$ (mod $p^{r}$). Then
$B_{r,\eta}=p^{2rm}+\sum_{j=0\atop\text{$r-j$ even}}^{r-1}p^{rm+jm-r}(p^{r-j}-p^{r-j-1})p^{\frac{r+j}{2}}.$
If $k<r$,
$B_{r,\eta}=\sum_{j=0\atop\text{$r-j$ even}}^{k}p^{rm+jm+\frac{-r+j}{2}}C_{r,j}+\sum_{j=0\atop\text{$r-j$ odd}}^{k}p^{rm+jm+\frac{-r+j}{2}}D_{r,j},$
where $A_{j}=p^{-2j}q(\eta)$, and
$C_{r,j}=\begin{cases}p^{r-j}-p^{r-j-1},&\text{if $p^{r-j}|A_{j}$}\\ -p^{r-j-1},&\text{if $p^{r-j}\nmid A_{j}$ but $p^{r-j-1}|A_{j}$}\\ 0,&\text{if $p^{r-j-1}\nmid A_{j}$ ($r-j>1$)}\end{cases},$
$D_{r,j}=\begin{cases}0,&\text{if $p^{r-j}|A_{j}$}\\ \Big(\frac{q(\eta)_{1}}{p}\Big)p^{r-j-\frac{1}{2}},&\text{if $p^{r-j}\nmid A_{j}$ but $p^{r-j-1}|A_{j}$}\\ 0,&\text{if $p^{r-j-1}\nmid A_{j}$ ($r-j>1$)}\end{cases}.$
Proof of Theorem 4.3. Suppose $2k+k^{\prime}$ is odd. Consider $(1-p^{2(m-s)})^{-1}J(s,\eta,\Phi_{p})$. Then by change of variables,
$(1-p^{2(m-s)})^{-1}J(s,\eta,\Phi_{p})=\sum_{b=0}^{\infty}\left(\sum_{r=0\atop\text{$b-r$ even}}^{b}p^{-mr}B_{r,\eta}\right)p^{(m-s)b}.$
[…]
We write $C(k,k^{\prime},b)=S_{\leq k}+S_{\geq k+1}$ where
$S_{\leq k}:=\sum_{\begin{subarray}{c}0\leq r\leq k\\ r\equiv b(2)\end{subarray}}p^{-mr}B_{r,\eta},\quad S_{\geq k+1}:=\sum_{\begin{subarray}{c}k+1\leq r\leq b\\ r\equiv b(2)\end{subarray}}p^{-mr}B_{r,\eta}.$
We first evaluate $S_{\leq k}$. From the definition of $B_{r,\eta}$ for $r\leq k$ we have
$S_{\leq k}=\sum_{\begin{subarray}{c}0\leq r\leq k\\ r\equiv b(2)\end{subarray}}p^{rm}+\sum_{\begin{subarray}{c}0\leq r\leq k\\ r\equiv b(2)\end{subarray}}\;\sum_{\begin{subarray}{c}0\leq j\leq r-1\\ r-j\ \mathrm{even}\end{subarray}}p^{jm-r}\Bigl(p^{\,r-j}-p^{\,r-j-1}\Bigr)p^{\frac{r+j}{2}}.$
Write $r=j+2t$ with $t\geq 1$. Then necessarily $r\equiv j\ (\mathrm{mod}\ 2)$, hence $j\equiv b\ (\mathrm{mod}\ 2)$. As for the second term of $S_{\leq k}$,
$p^{jm-r}\,p^{\frac{r+j}{2}}\Bigl(p^{\,r-j}-p^{\,r-j-1}\Bigr)=p^{jm+t}-p^{jm+t-1}.$
Therefore, for each fixed $j$ with $0\leq j\leq k$ and $j\equiv b(2)$,
$\sum_{t=1}^{\lfloor(k-j)/2\rfloor}\bigl(p^{jm+t}-p^{jm+t-1}\bigr)=p^{\,jm+\lfloor(k-j)/2\rfloor}-p^{jm}.$
Using $\sum_{\begin{subarray}{c}0\leq r\leq k\\ r\equiv b(2)\end{subarray}}p^{rm}=\sum_{\begin{subarray}{c}0\leq j\leq k\\ j\equiv b(2)\end{subarray}}p^{jm}$, we have
(11.7) $S_{\leq k}=\sum_{\begin{subarray}{c}0\leq j\leq k\\ j\equiv b(2)\end{subarray}}p^{\,jm+\lfloor(k-j)/2\rfloor}.$
Next we study $S_{\geq k+1}$. For $r\geq k+1$, by definition
$S_{\geq k+1}=\sum_{\begin{subarray}{c}k+1\leq r\leq b\\ r\equiv b(2)\end{subarray}}\;\sum_{\begin{subarray}{c}0\leq j\leq k\\ r-j\ \mathrm{even}\end{subarray}}p^{jm+\frac{-r+j}{2}}\;C_{r,j}.$
We write
$S_{\geq k+1}=U+V,\ \text{ where }\begin{cases}U:=\displaystyle\sum_{\begin{subarray}{c}k+1\leq r\leq b\\ r\equiv b(2)\end{subarray}}\ \sum_{\begin{subarray}{c}0\leq j\leq k\\ r-j\ \mathrm{even}\\ r+j\leq 2k+k^{\prime}\end{subarray}}p^{jm+\frac{-r+j}{2}}\bigl(p^{\,r-j}-p^{\,r-j-1}\bigr),\\[12.0pt] V:=\displaystyle\sum_{\begin{subarray}{c}k+1\leq r\leq b\\ r\equiv b(2)\end{subarray}}\ \sum_{\begin{subarray}{c}0\leq j\leq k\\ r-j\ \mathrm{even}\\ r+j=2k+k^{\prime}+1\end{subarray}}\Bigl(-\,p^{jm+\frac{-r+j}{2}}\,p^{\,r-j-1}\Bigr).\end{cases}$
Write $r=j+2t$ again. The parity condition $r\equiv b(2)$ yields $j\equiv b(2)$. The lower bound $r\geq k+1$ gives $t\geq t_{\min}(j):=\bigl\lceil(k+1-j)/2\bigr\rceil$. The inequality $r+j\leq 2k+k^{\prime}$ gives $t\leq t_{\max}(j):=\bigl\lfloor(2k+k^{\prime}-2j)/2\bigr\rfloor$. As for $U$, since $b\geq 2k+k^{\prime}$,
$p^{jm+\frac{-r+j}{2}}\bigl(p^{\,r-j}-p^{\,r-j-1}\bigr)=p^{jm+t}-p^{jm+t-1}.$
Thus
(11.8) $U=\sum_{\begin{subarray}{c}0\leq j\leq k\\ j\equiv b(2)\end{subarray}}\ \sum_{t=t_{\min}(j)}^{t_{\max}(j)}\bigl(p^{jm+t}-p^{jm+t-1}\bigr)=\sum_{\begin{subarray}{c}0\leq j\leq k\\ j\equiv b(2)\end{subarray}}\Bigl(p^{jm+t_{\max}(j)}-p^{jm+t_{\min}(j)-1}\Bigr).$
As for $V$, the condition $r+j=2k+k^{\prime}+1$ yields $t=t_{\max}(j)+1$. Since $b\geq 2k+k^{\prime}$,
(11.9) $V=\sum_{\begin{subarray}{c}0\leq j\leq k\\ j\equiv b(2)\end{subarray}}\Bigl(-\,p^{jm+t_{\max}(j)}\Bigr).$
Combining (11.8) and (11.9) yields
(11.10) $S_{\geq k+1}=-\sum_{\begin{subarray}{c}0\leq j\leq k\\ j\equiv b(2)\end{subarray}}p^{\,jm+t_{\min}(j)-1},\qquad t_{\min}(j)=\Bigl\lceil\frac{k+1-j}{2}\Bigr\rceil.$
For every integer $0\leq j\leq k$,
$\Bigl\lfloor\frac{k-j}{2}\Bigr\rfloor=\Bigl\lceil\frac{k+1-j}{2}\Bigr\rceil-1.$
Using (11.7) and (11.10),
$S_{\leq k}=\sum_{\begin{subarray}{c}0\leq j\leq k\\ j\equiv b(2)\end{subarray}}p^{\,jm+\lfloor(k-j)/2\rfloor}=\sum_{\begin{subarray}{c}0\leq j\leq k\\ j\equiv b(2)\end{subarray}}p^{\,jm+t_{\min}(j)-1}=-\,S_{\geq k+1}.$
Hence $C(k,k^{\prime},b)=S_{\leq k}+S_{\geq k+1}=0$.
For the functional equation, we compute the coefficients $C(k,k^{\prime},b)$ explicitly case by case. We compute one case, $k$ even. Let $k=2a$ and $2k+k^{\prime}=2l+1$. Then we prove
[…]
Suppose $b$ is odd. Let $b=2j+1$. Then if $b\leq k$,
$\sum_{r=0\atop\text{$r$ odd}}^{b}B_{r,\eta}p^{-rm}=\sum_{i=0}^{j}B_{2i+1,\eta}p^{-m(2i+1)}$
$=\sum_{i=0}^{j}(p^{m(2i+1)}+\sum_{u=0}^{i-1}p^{m(2u+1)}(p^{i-u}-p^{i-u-1}))=\sum_{u=0}^{j}p^{m(2u+1)}p^{j-u}.$
If $k\leq b<k+k^{\prime}$,
$\sum_{r=0\atop\text{$r$ odd}}^{b}B_{r,\eta}p^{-rm}=\sum_{i=0}^{a-1}+\sum_{i=a}^{j}.$
Here by change of variables,
$\sum_{i=0}^{a-1}=\sum_{u=0}^{a-1}p^{m(2u+1)}p^{a-1-u}.$
$\sum_{i=a}^{j}=\sum_{i=a}^{j}\sum_{u=0\atop i+u\leq l-1}^{a-1}p^{m(2u+1)}(p^{i-u}-p^{i-u-1})=\sum_{u=0}^{a-1}p^{m(2u+1)}(p^{j-u}-p^{a-u-1}),$
since $i+u\leq j+a-1\leq l-1$. So $i+u=l$ cannot occur. Hence if $k\leq b<k+k^{\prime}$,
$\sum_{r=0\atop\text{$r$ odd}}^{b}B_{r,\eta}p^{-rm}=\sum_{u=0}^{a-1}p^{m(2u+1)}p^{j-u}.$
If $b\geq k+k^{\prime}$,
$\sum_{r=0\atop\text{$r$ odd}}^{b}B_{r,\eta}p^{-rm}=\sum_{i=0}^{a-1}+\sum_{i=a}^{j},$
where
$\sum_{i=0}^{a-1}=\sum_{u=0}^{a-1}p^{m(2u+1)}p^{a-1-u},$
$\sum_{i=a}^{j}=\sum_{i=a}^{j}\sum_{u=0\atop i+u\leq l-1}^{a-1}p^{m(2u+1)}(p^{i-u}-p^{i-u-1})+\sum_{i=a}^{j}\sum_{u=0\atop i+u=l}^{a-1}p^{m(2u+1)}(-p^{i-u-1}).$
Now the first sum is
$\sum_{u=0}^{a-1}\sum_{i=a}^{\text{min$(j,l-1-u)$}}p^{m(2u+1)}(p^{i-u}-p^{i-u-1}))$
$=\sum_{u=0}^{l-1-j}p^{m(2u+1)}p^{j-u}+\sum_{u=l-j}^{a-1}p^{m(2u+1)}p^{l-2u-1}-\sum_{u=0}^{a-1}p^{m(2u+1)}p^{a-u-1}.$
In the second sum, if j<luj<l-u, u<lju<l-j, and it is an empty sum. If $j\geq l-u$, $u\geq l-j$, and it becomes $\sum_{u=l-j}^{a-1}p^{m(2u+1)}(-p^{l-2u-1})$. By summing up, we have, if $b\geq k+k^{\prime}$,
$\sum_{r=0\atop\text{$r$ odd}}^{b}B_{r,\eta}p^{-rm}=\sum_{u=0}^{l-1-j}p^{m(2u+1)}p^{j-u}.$
The case of $b$ even is similar.
Suppose $2k+k^{\prime}$ is even. Let $\varepsilon=\chi_{\eta}(p)$, and consider $(1+\varepsilon p^{m-s})^{-1}J(s,\eta,\Phi_{p})$. Then by change of variables,
$(1+\varepsilon p^{m-s})^{-1}J(s,\eta,\Phi_{p})=\sum_{b=0}^{\infty}(-\varepsilon)^{b}p^{(m-s)b}\left(\sum_{r=0}^{b}(-\varepsilon)^{r}p^{-mr}B_{r,\eta}\right).$
Let $C^{\varepsilon}(k,k^{\prime},b)=\sum_{r=0}^{b}(-\varepsilon)^{r}p^{-mr}B_{r,\eta}$. We prove that $C^{\varepsilon}(k,k^{\prime},b)=0$ if $b\geq 2k+k^{\prime}+1$.
We write $C^{\varepsilon}(k,k^{\prime},b)=S_{\leq k}+S_{\geq k+1}$ where
$S_{\leq k}:=\sum_{0\leq r\leq k}(-\varepsilon)^{r}p^{-mr}B_{r,\eta},\quad S_{\geq k+1}:=\sum_{k+1\leq r\leq b}(-\varepsilon)^{r}p^{-mr}B_{r,\eta}.$
We first evaluate $S_{\leq k}$:
$S_{\leq k}=\sum_{0\leq r\leq k}(-\varepsilon)^{r}p^{rm}+\sum_{0\leq r\leq k}\;\sum_{\begin{subarray}{c}0\leq j\leq r-1\\ r-j\ \mathrm{even}\end{subarray}}(-\varepsilon)^{r}\,p^{jm-r}\Bigl(p^{\,r-j}-p^{\,r-j-1}\Bigr)p^{\frac{r+j}{2}}.$
Put $r=j+2t$ with $t\geq 1$. Then $r\equiv j\ (\mathrm{mod}\ 2)$ and $(-\varepsilon)^{r}=(-\varepsilon)^{j}$. Since
$p^{jm-r}\,p^{\frac{r+j}{2}}\Bigl(p^{\,r-j}-p^{\,r-j-1}\Bigr)=p^{jm+t}-p^{jm+t-1}.$
for each fixed $j$ with $0\leq j\leq k$,
$\sum_{t=1}^{\lfloor(k-j)/2\rfloor}\bigl(p^{jm+t}-p^{jm+t-1}\bigr)=p^{\,jm+\lfloor(k-j)/2\rfloor}-p^{jm}.$
Using $\sum_{0\leq r\leq k}(-\varepsilon)^{r}p^{rm}=\sum_{0\leq j\leq k}(-\varepsilon)^{j}p^{jm}$, we obtain
$S_{\leq k}=\sum_{0\leq j\leq k}(-\varepsilon)^{j}\,p^{\,jm+\lfloor(k-j)/2\rfloor}.$
Next we evaluate $S_{\geq k+1}$. For $r\geq k+1$,
$S_{\geq k+1}=\sum_{k+1\leq r\leq b}\;\sum_{\begin{subarray}{c}0\leq j\leq k\\ r-j\ \mathrm{even}\end{subarray}}(-\varepsilon)^{r}\,p^{jm+\frac{-r+j}{2}}\,C_{r,j}+\sum_{k+1\leq r\leq b}\;\sum_{\begin{subarray}{c}0\leq j\leq k\\ r-j\ \mathrm{odd}\end{subarray}}(-\varepsilon)^{r}\,p^{jm+\frac{-r+j}{2}}\,D_{r,j}.$
Write $r=j+2t$ in the first double sum and $r=j+2t+1$ in the second. For the first sum, the inequalities $r\geq k+1$ and $r+j\leq 2k+k^{\prime}$ are equivalent to
$t_{\min}(j)\leq t\leq t_{\max}(j),\qquad t_{\min}(j):=\Bigl\lceil\frac{k+1-j}{2}\Bigr\rceil,\quad t_{\max}(j):=\Bigl\lfloor\frac{2k+k^{\prime}-2j}{2}\Bigr\rfloor.$
On this range,
$(-\varepsilon)^{r}\,p^{jm+\frac{-r+j}{2}}\,C_{r,j}=(-\varepsilon)^{j}\bigl(p^{jm+t}-p^{jm+t-1}\bigr).$
Hence for a fixed $j$,
$\sum_{t=t_{\min}(j)}^{t_{\max}(j)}(-\varepsilon)^{r}\,p^{jm+\frac{-r+j}{2}}\,C_{r,j}=(-\varepsilon)^{j}\Bigl(p^{jm+t_{\max}(j)}-p^{jm+t_{\min}(j)-1}\Bigr).$
For the second sum (odd rjr-j), nonzero terms occur only when $r+j=2k+k^{\prime}+1$ and it corresponds to $t=t_{\max}(j)$. Using $(-\varepsilon)^{r}\varepsilon=-(-\varepsilon)^{j}$, we obtain
$\sum_{\begin{subarray}{c}r=j+2t+1\\ r+j=2k+k^{\prime}+1\end{subarray}}(-\varepsilon)^{r}\,p^{jm+\frac{-r+j}{2}}\,D_{r,j}=-\,(-\varepsilon)^{j}\,p^{\,jm+t_{\max}(j)}.$
Since $b>2k+k^{\prime}$, summing over all $j$ yields
$S_{\geq k+1}=-\sum_{0\leq j\leq k}(-\varepsilon)^{j}\,p^{\,jm+t_{\min}(j)-1}.$
For every integer $0\leq j\leq k$,
$\Bigl\lfloor\frac{k-j}{2}\Bigr\rfloor=\Bigl\lceil\frac{k+1-j}{2}\Bigr\rceil-1=t_{\min}(j)-1.$
Hence
$S_{\leq k}=\sum_{0\leq j\leq k}(-\varepsilon)^{j}\,p^{\,jm+\lfloor(k-j)/2\rfloor}=\sum_{0\leq j\leq k}(-\varepsilon)^{j}\,p^{\,jm+t_{\min}(j)-1}=-\,S_{\geq k+1}.$
Therefore, $C^{\varepsilon}(k,k^{\prime},b)=S_{\leq k}+S_{\geq k+1}=0$.
For the functional equation, let $2k+k^{\prime}=2l$, and we compute explicitly $C^{\varepsilon}(k,k^{\prime},b)$: For $b<k$,
(11.12) $C^{\varepsilon}(k,k^{\prime},b)=\sum_{0\leq j\leq b}(-\varepsilon)^{j}\,p^{\,jm+\lfloor(b-j)/2\rfloor}.$
For $k\leq b\leq k+k^{\prime}$, $r+j\leq 2l$. Hence in this range, there is no contribution from $D_{r,j}=0$, and
$S_{\geq k+1}=\sum_{k+1\leq r\leq b}\;\sum_{\begin{subarray}{c}0\leq j\leq k\\ r-j\ \mathrm{even}\end{subarray}}(-\varepsilon)^{r}\,p^{jm}(p^{\frac{r-j}{2}}-p^{\frac{r-j}{2}-1})$
$\phantom{xxxxx}=\sum_{j=0}^{k}(-\varepsilon)^{j}p^{jm}(p^{\lfloor\frac{b-j}{2}\rfloor}-p^{\lceil\frac{k+1-j}{2}\rceil-1})=-S_{\leq k}+\sum_{j=0}^{k}(-\varepsilon)^{j}p^{jm+\lfloor\frac{b-j}{2}\rfloor}.$
Hence for $k\leq b\leq k+k^{\prime}$, $C^{\varepsilon}(k,k^{\prime},b)=\sum_{j=0}^{k}(-\varepsilon)^{j}p^{jm+\lfloor\frac{b-j}{2}\rfloor}.$ For $0\leq i\leq k^{\prime}$, since $k^{\prime}$ is even, $\lfloor\frac{k+k^{\prime}-i-j}{2}\rfloor=\lfloor\frac{k+i-j}{2}\rfloor+\frac{k^{\prime}}{2}-i$. Hence
$C^{\varepsilon}(k,k^{\prime},k+k^{\prime}-i)=p^{k+k^{\prime}-i-l}C^{\varepsilon}(k,k^{\prime},k+i).$
For $k+k^{\prime}<b\leq 2k+k^{\prime}$,
$S_{\geq k+1}=\sum_{k+1\leq r\leq b}\;\sum_{\begin{subarray}{c}0\leq j\leq k\\ r-j\ \mathrm{even}\end{subarray}\atop\text{$r+j\leq 2l$}}(-\varepsilon)^{r}\,p^{jm}(p^{\frac{r-j}{2}}-p^{\frac{r-j}{2}-1})+\sum_{k+1\leq r\leq b}\;\sum_{\begin{subarray}{c}0\leq j\leq k\\ \text{$r+j=2l+1$}\end{subarray}}(-\varepsilon)^{r}\,p^{jm}(\varepsilon p^{\frac{r-j-1}{2}}).$
The second sum is $\sum_{j=2l-b+1}^{k}(-\varepsilon)^{j}p^{jm+l-j}.$ The first sum is
$\sum_{j=0}^{2l-b}(-\varepsilon)^{j}p^{jm}\sum_{r=k+1}^{b}(p^{\frac{r-j}{2}}-p^{\frac{r-j}{2}-1})+\sum_{j=2l-b+1}^{k}(-\varepsilon)^{j}p^{jm}\sum_{r=k+1}^{2l-j}(p^{\frac{r-j}{2}}-p^{\frac{r-j}{2}-1})$
$\phantom{xxxxxxxxxxxxx}=-S_{\leq k}+\sum_{j=0}^{2l-b}(-\varepsilon)^{j}p^{\lfloor\frac{b-j}{2}\rfloor}+\sum_{j=2l-b+1}^{k}(-\varepsilon)^{j}p^{jm+l-j}.$
Therefore, if $k+k^{\prime}<b\leq 2k+k^{\prime}$,
$C^{\varepsilon}(k,k^{\prime},b)=\sum_{j=0}^{2l-b}(-\varepsilon)^{j}p^{jm+\lfloor\frac{b-j}{2}\rfloor}.$
Hence, for $0\leq i<k$, $C^{\varepsilon}(k,k^{\prime},2k+k^{\prime}-i)=p^{l-i}C^{\varepsilon}(k,k^{\prime},i).$ This proves the functional equation. $\square$
11.2.2. The case when $p=2$ (and $n$ is odd)
When $p=2$, we need the analogue of (11.6), the following generalized Gauss sums.
Lemma 11.2.
For $r\geq 1$, and $x,a\in\mathbb{Z}$, define
$G(x,a):=\sum_{u\ (\mathrm{mod}\ 2^{r})}\exp\!\Bigl(\frac{2\pi i}{2^{r}}\bigl(xu^{2}+2au\bigr)\Bigr).$
When $x\neq 0$, write $x=2^{j}x_{0}$ with $2\nmid x_{0}$.
$G(x,a)=\begin{cases}2^{r},&j\geq r,\ 2^{r-1}\mid a,\,r\geq 1,\\[2.0pt] 2^{r},&j=r-1,\ 2^{j}\|2a,\\[2.0pt] 2^{\frac{r+j}{2}}\exp\!\Bigl(-\frac{2\pi i}{2^{r-j}}\cdot(a^{\prime})^{2}x_{0}^{-1}\Bigr)\,\Big(\frac{2}{x_{0}}\Big)^{r-j}(1+i^{\,x_{0}}),&j\leq r-2,\,2^{j+1}|2a,\\ 0,&\text{otherwise}\end{cases}$
where $(\frac{2}{x_{0}})$ is the Kronecker symbol, and $x_{0}^{-1}$ is taken modulo $2^{r-j}$, and $a=2^{j}a^{\prime}$.
Proof.
If $j\geq r$, then $x\equiv 0\ (\mathrm{mod}\ 2^{r})$ and the claim is easy to check. Assume $0\leq j<r$. Write $x=2^{j}x_{0}$ with $2\nmid x_{0}$, we write
$u=u_{0}+2^{\,r-j}t,\qquad u_{0}\ (\mathrm{mod}\ 2^{\,r-j}),\quad t=0,1,\dots,2^{j}-1.$
Then we see that $xu^{2}+2au\equiv 2^{j}x_{0}u_{0}^{2}+2au_{0}+2a2^{\,r-j}t\pmod{2^{r}}$. Therefore
$G(x,a)=\sum_{u_{0}\ (\mathrm{mod}\ 2^{\,r-j})}\exp\!\Bigl(\frac{2\pi i}{2^{\,r-j}}\bigl(x_{0}u_{0}^{2}+(2a/2^{j})u_{0}\bigr)\Bigr)\cdot\sum_{t=0}^{2^{j}-1}\exp\!\Bigl(\frac{2\pi i}{2^{j}}\,2a\,t\Bigr).$
The latter sum in $t$ equals 0 unless $2^{j}\mid 2a$, and equals $2^{j}$ when $2a=2^{j}a^{\prime\prime}$. Thus, if $2^{j}\nmid 2a$ then $G(x,a)=0$. If $2a=2^{j}a^{\prime\prime}$, then
$G(x,a)=2^{j}\sum_{u_{0}\ (\mathrm{mod}\ 2^{\,r-j})}\exp\!\Bigl(\frac{2\pi i}{2^{\,r-j}}\bigl(x_{0}u_{0}^{2}+a^{\prime\prime}u_{0}\bigr)\Bigr).$
Put T:=rjT:=r-j and $a^{\prime\prime}:=a/2^{\,j}$. Then
$G(x,a)=2^{j}\,S_{T}(x_{0},a^{\prime\prime}),\qquad S_{T}(x_{0},a^{\prime\prime}):=\sum_{u_{0}\ (\mathrm{mod}\ 2^{T})}\exp\!\Bigl(\frac{2\pi i}{2^{T}}\,\bigl(x_{0}u_{0}^{2}+a^{\prime\prime}u_{0}\bigr)\Bigr).$
If $T=1$, a direct check gives $S_{1}(x_{0},a^{\prime\prime})=2$ for $a^{\prime\prime}$ odd and 0 for $a^{\prime\prime}$ even. If $T\geq 2$ and $a^{\prime\prime}$ is odd, we pair $u_{0}$ with $u_{0}+2^{T-1}$ to see cancellation so that $S_{T}(x_{0},a^{\prime\prime})=0$. If $T\geq 2$ and $a^{\prime\prime}$ is even (say $a^{\prime\prime}=2a^{\prime}$), we have $x_{0}u_{0}^{2}+a^{\prime\prime}u_{0}=x_{0}(u_{0}+a^{\prime}x_{0}^{-1})^{2}-x_{0}(a^{\prime}x_{0}^{-1})^{2},$ where $x_{0}^{-1}$ is taken modulo $2^{T}$. Hence
$S_{T}(x_{0},a^{\prime\prime})=\exp\!\Bigl(-\frac{2\pi i}{2^{T}}\,(a^{\prime})^{2}x_{0}^{-1}\Bigr)\sum_{v\ (\mathrm{mod}\ 2^{T})}\exp\!\Bigl(\frac{2\pi i}{2^{T}}\,x_{0}v^{2}\Bigr)=\exp\!\Bigl(-\frac{2\pi i}{2^{T}}\cdot(a^{\prime})^{2}x_{0}^{-1}\Bigr)\,\Big(\frac{2}{x_{0}}\Big)^{T}(1+i^{\,x_{0}})\,2^{T/2},$
where the final identity uses a well-known formula for quadratic Gauss sums (cf. [2, Theorem 1.5.4, p.27]). ∎
We also need the following [2, p. 29]:
Lemma 11.3.
Let $\chi$ be a nontrivial Dirichlet character mod $k$, and conductor $f$. Let $\tau_{k}(m,\chi)=\sum_{n=1}^{k-1}\chi(n)e^{2\pi imn/k}$. Then
$\tau_{k}(m,\chi)=\begin{cases}kf^{-1}\tau_{f}(fm/k,\chi),&\text{if $k|fm$}\\ 0,&\text{otherwise}\end{cases}.$
Moreover, $\tau_{f}(m,\chi)=\overline{\chi}(m)\tau_{f}(1,\chi)$. We have the following: $\tau_{4}(1,\chi)=2i$ if $\chi$ is non-trivial. Let $\chi$ be a primitive character (mod 8) normalized so that $\chi(5)=-1$. Then $\tau_{8}(1,\chi)=2\sqrt{2}i^{\frac{1-\chi(-1)}{2}}$.
Then
Proposition 11.4.
Suppose $T\geq 4$ is even and $m>0$. Then
$\sum_{x\in\mathbb{Z}/2^{T}\mathbb{Z}}\exp\left(-\frac{2\pi i}{2^{T}}xm\right)(\tfrac{2}{x})^{T}(1+i^{x})=\begin{cases}2^{T}-2^{T-1},&\text{if $2^{T}|m$}\\ -2^{T-1},&\text{if $2^{T-1}\|m$}\\ 2^{T-1}(\frac{-1}{m2^{2-T}}),&\text{if $2^{T-2}\|m$}\\ 0,&\text{otherwise}.\end{cases}$
When $T=2$, we have
$\sum_{x\in\mathbb{Z}/4\mathbb{Z}}\exp\left(-\frac{2\pi i}{4}xm\right)(\tfrac{2}{x})^{2}(1+i^{x})=\begin{cases}2,&\text{if $4|m$}\\ -2,&\text{if $2\|m$}\\ 2(\frac{-1}{m}),&\text{if $2\nmid m$}.\end{cases}$
Suppose $T\geq 5$ is odd. Then
$\sum_{x\in\mathbb{Z}/2^{T}\mathbb{Z}}\exp\left(-\frac{2\pi i}{2^{T}}xm\right)(\tfrac{2}{x})^{T}(1+i^{x})=\begin{cases}2^{T-\frac{3}{2}}(\frac{2}{m2^{3-T}})(1+(\frac{-1}{m2^{3-T}})),&\text{if $2^{T-3}\|m$}\\ 0,&\text{otherwise}.\end{cases}$
When $T=3$, we have
$\sum_{x\in\mathbb{Z}/8\mathbb{Z}}\exp\left(-\frac{2\pi i}{8}xm\right)(\tfrac{2}{x})^{3}(1+i^{x})=\begin{cases}2^{\frac{3}{2}}(\frac{2}{m})(1+(\frac{-1}{m})),&\text{if $2\nmid m$}\\ 0,&\text{otherwise}.\end{cases}$
Proof.
We only prove when $T$ is odd. Notice the sum runs over odd $x$ since $(\tfrac{2}{x})=0$ for $x$ even. Since $i^{x}=(\frac{-1}{x})i$ for $x$ odd, write the sum as
$\sum_{x\in\mathbb{Z}/2^{T}\mathbb{Z}}e^{-\frac{2\pi i}{2^{T}}xm}(\tfrac{2}{x})+i\sum_{x\in\mathbb{Z}/2^{T}\mathbb{Z}}e^{-\frac{2\pi i}{2^{T}}xm}(\tfrac{-2}{x}).$
Let $\chi(x)=(\tfrac{2}{x})$ and $\chi^{\prime}(x)=(\tfrac{-2}{x})$. They are primitive characters mod 8, and $\chi(5)=-1,\chi^{\prime}(5)=-1$. Then apply the above lemma. ∎
Now suppose $2^{k}\|\eta$ and $2^{2k+k^{\prime}}\|q(\eta)$. If $r\leq k$, then
$B_{r,\eta}=2^{2mr}+\sum_{j=0\atop\text{$r-j$ even}}^{r-1}2^{mr+mj+\frac{-r+j}{2}}(2^{r-j}-2^{r-j-1}).$
If $r>k$, we need to divide into two cases:
Case 1. $2^{k+1}|a_{1},...,a_{m},b_{1},...,b_{m}$, but $2^{k}\|a_{0};$ In this case, by congruence, $k^{\prime}=0$. If $r\geq k+3$,
$B_{r,\eta}=\sum_{j=0\atop\text{$r-j$ even}}^{k}2^{rm+jm+\frac{-r+j}{2}}C^{(2)}_{r,j}+\sum_{j=0\atop\text{$r-j$ odd}}^{k}2^{rm+jm+\frac{-r+j}{2}}D^{(2)}_{r,j}$,
where $A_{j}=2^{-2j}q(\eta)$, and [Note that since $j\leq k$, $r-j\geq 3$.]
$C^{(2)}_{r,j}=\begin{cases}2^{r-j}-2^{r-j-1},&\text{if $2^{r-j}|A_{j}$}\\ -2^{r-j-1},&\text{if $2^{r-j-1}\|A_{j}$}\\ 2^{r-j-1}(\frac{-1}{A_{j}2^{2-r+j}}),&\text{if $2^{r-j-2}\|A_{j}$}\\ 0,&\text{if $2^{r-j-2}\nmid A_{j}$}\end{cases},$
$D^{(2)}_{r,j}=\begin{cases}2^{r-j-\frac{3}{2}}(\tfrac{2}{A_{j}2^{3-r+j}})(1+(\tfrac{-1}{A_{j}2^{3-r+j}})),&\text{if $2^{r-j-3}\|A_{j}$ and $r-j\geq 3$}\\ 0,&\text{otherwise.}\end{cases}$
If $r=k+1$, there is no contribution from $j=k$ by Lemma 11.2. But there is contribution from $x=0$. [We put the convention that the empty sum is zero.] So
$B_{k+1,\eta}=2^{2m(k+1)}+\sum_{j=0\atop\text{$k+1-j$ even}}^{k}2^{(k+1)m+jm+\frac{-k-1+j}{2}}(2^{k+1-j}-2^{k-j}).$
If $r=k+2$, there is a contribution from $j=k+1$. Here we use the fact that $2^{k+1}|a_{1},...,b_{m}$.
$B_{k+2,\eta}=2^{(2k+3)m}+\sum_{j=0\atop\text{$k+2-j$ even}}^{k}2^{(k+2)m+jm+\frac{-k-2+j}{2}}(2^{k+2-j}-2^{k+1-j}).$
Case 2. $2^{k}\|(a_{1},...,a_{m},b_{1},...,b_{m})$ and $2^{k}|a_{0}$. If $r\geq k+2$,
$B_{r,\eta}=\sum_{j=0\atop\text{$r-j$ even}}^{k}2^{rm+jm+\frac{-r+j}{2}}C^{(2)}_{r,j}+\sum_{j=0\atop\text{$r-j$ odd}}^{k}2^{rm+jm+\frac{-r+j}{2}}D^{(2)}_{r,j}$.
If $r=k+1$, there is no contribution from $j=k$ term by Lemma 11.2. So
$B_{k+1,\eta}=\sum_{j=0\atop\text{$k+1-j$ even}}^{k}2^{(k+1)m+jm+\frac{-k-1+j}{2}}(2^{k+1-j}-2^{k-j}).$
Proof of Theorem 4.4. We first look at the special case of $k=0$.
Case 1. Then $k^{\prime}=0$, and
$J(s,\eta,\Phi_{2})=1+2^{2m}2^{-s}+(2^{3m}+2^{2m})2^{-2s}+2^{3m}(\tfrac{2}{q(\eta)})(1+(\tfrac{-1}{q(\eta)}))2^{-3s}.$
Now if $a_{0}$ is odd, $a_{0}^{2}\equiv 1$ (mod 8). Hence $q(\eta)\equiv 1$ (mod 8). Therefore, if $X=2^{m-s}$, $J(s,\eta,\Phi_{2})=(1+X)(1+(2^{m}-1)X+2X^{2})$. So $\chi_{\eta}(2)=1$, and $Q_{\eta}(x)=1+(2^{m}-1)X+2X^{2}$.
Case 2.
$J(s,\eta,\Phi_{2})=1+\sum_{r=2\atop\text{$r$ even}}^{k^{\prime}}2^{rm+\frac{r}{2}-1-rs}-\delta_{2,k^{\prime}+1}2^{m(k^{\prime}+1)+\frac{k^{\prime}+1}{2}-1-(k^{\prime}+1)s}$
$+\delta_{2,k^{\prime}+2}2^{m(k^{\prime}+2)+\frac{k^{\prime}}{2}-(k^{\prime}+2)s}(\tfrac{-1}{q(\eta)2^{-k^{\prime}}})+\delta_{1,k^{\prime}+3}2^{m(k^{\prime}+3)+\frac{k^{\prime}}{2}-(k^{\prime}+3)s}(\tfrac{2}{q(\eta)2^{-k^{\prime}}})(1+(\tfrac{-1}{q(\eta)2^{-k^{\prime}}})),$
where $\delta_{2,n}=\begin{cases}1,&\text{if $n$ is even}\\ 0,&\text{if $n$ is odd}\end{cases}$, and $\delta_{1,n}=\begin{cases}1,&\text{if $n$ is odd}\\ 0,&\text{if $n$ is even}\end{cases}$.
When $k^{\prime}$ is odd, let $k^{\prime}=2l+1$. Then
$J(s,\eta,\Phi_{2})=1+\sum_{a=1}^{l}2^{a-1}X^{a}-2^{l}X^{2l+2}=(1-X^{2})(1+\sum_{a=1}^{l}2^{a}X^{2a}).$
Since $\chi_{\eta}(a)=0$, $Q_{\eta}(X)=1+\sum_{a=1}^{l}2^{a}X^{2a}$.
When $k^{\prime}$ is even, and $q(\eta)2^{-k^{\prime}}\equiv 3$ (mod 4), it is similar to Case 1. Suppose $q(\eta)2^{-k^{\prime}}\equiv 1$ (mod 8). Let $k^{\prime}=2l$. Then
$J(s,\eta,\Phi_{2})=(1+X)(1+\sum_{a=1}^{l+1}2^{a}X^{2a}-\sum_{a=1}^{l+1}2^{a-1}X^{2a-1}).$
Since $\chi_{\eta}(2)=1$, $Q_{\eta}(X)=1+\sum_{a=1}^{l+1}2^{a}X^{2a}-\sum_{a=1}^{l+1}2^{a-1}X^{2a-1}$.
Suppose $q(\eta)2^{-k^{\prime}}\equiv 5$ (mod 8). In this case,
$J(s,\eta,\Phi_{2})=(1-X)(1+\sum_{a=1}^{l+1}2^{a}X^{2a}+\sum_{a=1}^{l+1}2^{a-1}X^{2a-1}).$
So $\chi_{\eta}(2)=-1$, $Q_{\eta}(X)=1+\sum_{a=1}^{l+1}2^{a}X^{2a}+\sum_{a=1}^{l+1}2^{a-1}X^{2a-1}$.
Now for a general $k>0$, by the definition of $C_{r,j}^{(2)}$ and $D_{r,j}^{(2)}$, we see that if $r>2k+k^{\prime}+2$, $C_{r,j}^{(2)}=0$ for any $j$, and if $r>2k+k^{\prime}+3$, $D_{r,j}^{(2)}=0$ for any $j$. So $J(s,\eta,\Phi_{2})$ is a polynomial in $2^{m-s}$.
If $k^{\prime}$ is odd, then $\chi_{\eta}(2)=0$, and it is exactly the same as odd prime case.
If $k^{\prime}$ is even, and $q(\eta)2^{-2k-k^{\prime}}\equiv 3$ (mod 4), then $\chi_{\eta}(2)=0$, and it is similar to odd $k^{\prime}$ case.
When $k^{\prime}$ is even, and $q(\eta)2^{-2k-k^{\prime}}\equiv 1$ (mod 8), then $\chi_{\eta}(2)=1$. In this case,
$Q_{\eta}(2^{m-s})=(1+2^{m-s})^{-1}J(s,\eta,\Phi_{2})=\sum_{b=0}^{\infty}(-1)^{b}2^{(m-s)b}C(k,k^{\prime},b),$
where $C(k,k^{\prime},b)=\sum_{r=0}^{b}(-1)^{r}2^{-mr}B_{r,\eta}$. Let $2k+k^{\prime}=2l$. We can compute the coefficients explicitly as the odd prime case, and show that $C(k,k^{\prime},2l+2-i)=2^{l+1-i}C(k,k^{\prime},i)$ for $0\leq i\leq k+1$. If $k+2\leq i\leq k+k^{\prime}$,
$C(k,k^{\prime},2l+2-i)=2^{l+1-i}C(k,k^{\prime},i).$
Similarly for $k^{\prime}$ even and $q(\eta)2^{-2k-k^{\prime}}\equiv 5$ (mod 8), we compute the coefficients of $Q_{\eta}(2^{m-s})=(1-2^{m-s})^{-1}J(s,\eta,\Phi_{2})$. Since it is similar to the proof of Theorem 4.3, we omit the details here. For the details, we refer to [17]. $\square$
12. Appendix
In this section, we explicitly compute $C(S,r)=C(r,\sqrt{\beta}^{-1}\lambda)$ in (7.11) as a generalization of [24, Theorem A.1]: Let
$I_{v}(r,\mu,\lambda)=\int_{\mathbb{R}}t^{r}e^{-t^{2}}\Bigg(\frac{(t+\lambda\sqrt{-1})^{2}-\mu}{|(t+\lambda\sqrt{-1})^{2}-\mu|}\Bigg)^{v}K_{v}(|(t+\lambda\sqrt{-1})^{2}-\mu|)dt.$
We first consider the case $r$ even.
Theorem 12.1.
Suppose $\mu,\lambda>0,\ r\in\mathbb{Z}_{\geq 0}$ and $v\in\mathbb{Z}$. Then
$I_{v}(2r,\mu,\lambda)=(-1)^{v}C(2r,\lambda)e^{-\mu},$
where
$C(2r,\lambda)=\pi e^{\lambda^{2}}\Gamma(\tfrac{1}{2},\lambda^{2})e^{\lambda^{2}}p_{r}(\lambda^{2})+\pi e^{-\lambda^{2}}\lambda q_{r-1}(\lambda^{2}).$
Here $\Gamma(\frac{1}{2},x)=\int_{x}^{\infty}t^{-1/2}e^{-t^{2}}\,dt$ is the incomplete Gamma function, and $p_{r},q_{r-1}$ are polynomials of degree r,r1r,r-1 (resp.) with rational coefficients. If $r=0$, set $q_{-1}=0$.
Let $I_{v}=I_{v}(r,\mu,\lambda)$. Then by the change of variables, we see that $I_{-v}=I_{v}$ and $\overline{I_{v}}=I_{v}$. So $I_{v}$ is real. Also [24, Proposition A.2] says, for any $r$,
(12.1) $\partial_{\mu}I_{v}(r,\mu,\lambda)=\frac{1}{2}(I_{v+1}(r,\mu,\lambda)+I_{v-1}(r,\mu,\lambda)).$
Now we explain how to modify the proof of [24, Proposition A.2] to obtain $I_{0}(2r,\mu,\lambda)=C(2r,\lambda)e^{-\mu}.$ Then together with the recurrence relation above, we obtain the theorem.
Let $s=\zeta t^{\frac{1}{2}}+\zeta^{-1}t^{-\frac{1}{2}}$, where $\zeta=e^{\sqrt{-1}\frac{\pi}{4}}$. Then $s^{2}=2+\sqrt{-1}(t-t^{-1})$ and $s\bar{s}=t+t^{-1}$. Let $\alpha=\mu+\lambda^{2}$. As in the proof of [24, Proposition A.2], we have
$I_{0}(2r,\mu,\lambda)=e^{-\mu-\lambda^{2}}\int_{0}^{\infty}\int_{0}^{\infty}t^{2r}e^{-\frac{1}{2}s^{2}x^{2}-\lambda s\bar{s}x+\frac{1}{2}\alpha s^{2}}\,\frac{dt}{t}dx.$
Now
$\int_{0}^{\infty}t^{2r}e^{-\frac{1}{4\beta}x^{2}-\gamma x}\,dx=4^{r}\beta^{r}\left(\sqrt{\beta}\Gamma(r+\tfrac{1}{2}){}_{1}F_{1}(r+\tfrac{1}{2},\tfrac{1}{2};\beta\gamma^{2})-2\beta\gamma\Gamma(r+1){}_{1}F_{1}(r+1,\tfrac{3}{2};\beta\gamma^{2}\right).$
Apply the above formula with $\beta=(s\sqrt{2})^{-2},\gamma=\lambda s\bar{s}$. Then
$e^{\mu}I_{0}(2r,\mu,\lambda)=$
$e^{-\lambda^{2}}\int_{0}^{\infty}e^{\frac{1}{2}\alpha s^{2}}\frac{2^{\frac{n}{2}}}{s^{2r}}\left(\tfrac{1}{s\sqrt{2}}\Gamma(r+\tfrac{1}{2}){}_{1}F_{1}(r+\tfrac{1}{2},\tfrac{1}{2};\tfrac{1}{2}\lambda^{2}{\bar{s}}^{2})-\lambda\tfrac{\bar{s}}{s}\Gamma(1+r){}_{1}F_{1}(r+1,\tfrac{3}{2};\tfrac{1}{2}\lambda^{2}{\bar{s}}^{2})\right)\,\frac{dt}{t}.$
Let $u=-\sqrt{-1}\frac{\pi}{4}+\frac{1}{2}\log t$ so that $du=\frac{dt}{2t}$, and $\cosh(u)=\frac{\bar{s}}{2}$, and $\sinh(u)=-\frac{is}{2}$. Then
$e^{\mu}I_{0}(2r,\mu,\lambda)=e^{-\lambda^{2}}\int_{Im(u)=-\sqrt{-1}\frac{\pi}{2}}e^{-2\alpha(\sinh(u))^{2}}\frac{1}{(-2)^{r}(\sinh(u))^{2r}}$
$\cdot\left(\tfrac{1}{2i\sqrt{2}\sinh(u)}\Gamma(r+\tfrac{1}{2}){}_{1}F_{1}(r+\tfrac{1}{2},2\lambda^{2}(\cosh(u))^{2})-\lambda\tfrac{\cosh(u)}{i\sinh(u)}\Gamma(1+r){}_{1}F_{1}(r+1,\tfrac{3}{2};2\lambda^{2}(\cosh(u))^{2})\right)\,\frac{dt}{t}.$
We differentiate under the integral with respect to $\mu$ and move the contour to $Im(u)=0$. Then the integral becomes an integral of an odd function, and it is zero. Therefore, $I_{0}(2r,\mu,\lambda)=C(2r,\lambda)e^{-\mu}$ for some $C(2r,\lambda)$.
In order to compute $C(2r,\lambda)$, consider $\lim_{\mu\to 0}I_{0}(2r,\mu,\lambda)$. Then
$C(2r,\lambda)=\int_{-\infty}^{\infty}t^{2r}e^{-t^{2}}K_{0}(t^{2}+\lambda^{2})\,dt.$
We use the substitution $t^{2}+\lambda^{2}=u$. Then
$C(2r,\lambda)=e^{\lambda^{2}}\int_{\lambda^{2}}^{\infty}(u-\lambda^{2})^{r-\frac{1}{2}}e^{-u}K_{0}(u)\,du.$
We note that $C(0,\lambda)=\frac{\pi}{\sqrt{2}}\Gamma(\frac{1}{2},2\lambda^{2})$.
Let
$D(r,\lambda)=e^{\lambda^{2}}\int_{\lambda^{2}}^{\infty}(u-\lambda^{2})^{r-\frac{1}{2}}e^{-u}K_{1}(u)\,du.$
Then by using the fact that $\frac{dK_{0}(t)}{dt}=K_{1}(t)$, we can show
$C(2r,\lambda)+D(r,\lambda)=(r-\frac{1}{2})C(2r-2,\lambda),$
$\frac{d}{d\lambda}C(2r,\lambda)=-2\lambda D(r,\lambda).$
So we have the differential equation,
$\frac{d}{d\lambda}C(2r,\lambda)-2\lambda C(2r,\lambda)=2\lambda(r-\frac{1}{2})C(2r-2,\lambda),$
Using this recurrence relation, we can show the result.
Remark 12.2.
Suppose $r$ is odd. Then it is clear that $I_{0}(r,\mu,\lambda)=0$. By the change of variables, we can see that $I_{-v}=-I_{v}$ and $\overline{I_{v}}=I_{-v}$. Hence $I_{v}$ is purely imaginary. Let $z=(t+\lambda\sqrt{-1})^{2}-\mu$. Then
$I_{1}(r,\mu,\lambda)=4\lambda\sqrt{-1}\int_{0}^{\infty}t^{r+1}e^{-t^{2}}\frac{K_{1}(|z|)}{|z|}\,dt.$
For $v>1$, $I_{v}(r,\mu,\lambda)$ can be obtained from $I_{1}(r,\mu,\lambda)$ by (12.1), but we may not have a similar result as in the case when $r$ is even.